### v1

# On thermal transpiration and thermomolecular pressure difference

Abstract
In this article, we demonstrate the phenomenon of thermal transpiration in a bounded convex domain. We employ the stationary Boltzmann equation with a cutoff potential. For boundary condition, we partition the boundary into diffuse reflection and incoming regions. We establish the existence of solution in a weighted $L^{\infty}$ space. Furthermore, we consider a convex domain with diffuse reflection boundary condition in the middle and incoming boundary condition at the two ends. We first consider Maxwellians with the same pressure but different temperatures at the two ends. We prove that the total flux $U(x)$ is directed toward the hot end. Furthermore, we derive an estimate for the total flux:
$U(x)\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (1)
In addition, we show that when the pressures and temperatures on the two ends satisfy the relation
$\frac{P_{1}}{P_{2}}=\sqrt{\frac{T_{1}}{T_{2}}},$ (2)
the total flux of the solution is of order $\mathcal{O}(\frac{1}{\kappa})$. This result is consistent with Knudsen’s finding of thermomolecular pressure difference in 1909.
Chapter 1. Introduction
1.1. Introduction
The thermal transpiration is a phenomenon that temperature gradient on the boundary induces a flow in a pipe [36]. This phenomenon has many engineering applications, including the Crookes radiometer [11], which can be used in energy harvesting [37]; the Knudsen compressor, which pumps gas without any moving mechanical part [43]; microelectromechanical systems [32], and MEMS-based multistage Knudsen compressor, where the gas moves from a cold chamber to a hot chamber creating a pressure difference across the compressor element [21]. Further applications and details can be found in [23], [36].
The thermal transpiration was first observed by Reynolds in 1879 [35]. He observed that gas moves along a solid surface due to temperature differences. In the same year, Maxwell developed a theory to explain this effect [29]. Moreover, when considering two chambers connected by a tube, Knudsen showed that in the limit of infinitely small tube, the pressure between the two chambers satisfies
$\frac{P_{1}}{\sqrt{T_{1}}}=\frac{P_{2}}{\sqrt{T_{2}}}$ (1.1)
in equilibrium. This phenomenon is called the thermomolecular pressure difference. This relation is discussed in [25], [26]. In an experiment, S. Chu Liang provided a more accurate expression of (1.1) in [27], with further applications given in [2], [40].
In kinetic theory, the stationary Boltzmann equation serves as a fundamental mathematical model for describing the behavior of rarefied gas. It reads
$v\cdot\nabla_{x}F(x,v)=\frac{1}{\kappa}Q(F,F)(x,v),\quad\text{in}\quad\Omega\times\mathbb{R}^{3}.$ (1.2)
Here, $\kappa$ is Knudsen number and $F$ represents the velocity distribution function of the gas molecules. The velocity distribution function is used to describe the number of molecules $\mathrm{d}N$ in a six-dimensional phase-space element $\mathrm{d}\mathbf{X}\mathrm{d}\bm{\xi}$. That is,
$\mathrm{d}N=F(\mathbf{X},\bm{\xi})\,\mathrm{d}\mathbf{X}\mathrm{d}\bm{\xi}.$
The collision operator $Q$ that appears in (1.2) is defined as follows.
$Q(F,G)(x,v)$
$=\int_{\mathbb{R}^{3}}\int_{S^{2}}\left[F(v^{\prime})G(u^{\prime})+F(u^{\prime})G(v^{\prime})-F(v)G(u)-F(u)G(v)\right]B(|v-u|,\omega)d\omega du,$
where
$v^{\prime}=v+((u-v)\cdot\omega)\omega\quad\text{and}\quad u^{\prime}=u-((u-v)\cdot\omega)\omega.$ (1.3)
Note that $v^{\prime},u^{\prime}$ can also be expressed in the equivalent form
$v^{\prime}=\frac{v+u}{2}+\frac{|v-u|}{2}\sigma\quad\text{and}\quad u^{\prime}=\frac{v+u}{2}-\frac{|v-u|}{2}\sigma,$ (1.4)
where $\sigma\in\mathbb{S}^{2}$.
Consequently, stationary Boltzmann equation has been extensively applied to describe the gas distribution in the domains involving thermal transpiration. However, in engineering applications, obtaining numerical solutions requires substantial computational efforts, either deterministically ([38], [42], [45]) or stochastically ([3], [33]). Some numerical analysis of thermal transpiration employ the lattice Boltzmann model [39], the BGK model ([20], [28], [30]) or the linearized Boltzmann Equation model [41]. In particular, for the linearized Boltzmann equation model, Sone, Y. and Aoki, K. have made many significant contributions ([KazuoAokiYoshioSoneandYorifumiWaniguchi], [31]). Moreover, Sone provided an explicit form of the solution when the free transport equation is used to describe the thermal transpiration [36].
Although there are many results on the numerical analysis of thermal transpiration using the linearized Boltzmann equation, there are only a few mathematical results. In 2007, Chiun-Chuan Chen, I-Kun Chen, Tai-Ping Liu and Yoshio Sone solved this problem for gas between two plates or within an infinite long pipe under diffuse reflection boundary conditions. They showed that when the Knudsen number $\kappa$ is sufficiently large, the local flow between two plates was of the order log $\kappa$ and the flow in a pipe has finite positive velocity. Moreover, they proved that in both case, the local flux is proportional to the wall temperature gradient [6]. This result is used to study the boundary logarithmic singularity ([9],[41]).
In this paper, we address the problem in the case where the pipe has finite length, and we impose the incoming boundary condition at both ends with Maxwellians and diffusion boundary condition in the middle.
In this article, we formulate the mixed boundary problem for the Boltzmann equation as follows:
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=F_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ F(x,v)&=m_{T(x)}\int_{u\cdot n(x)>0}F(x,u)u\cdot n(x)du,&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (1.5)
where $\Omega$ is a bounded convex domain in $\mathbb{R}^{3}$. Here, $\partial\Omega_{1}$ and $\partial\Omega_{2}$ are subsets of $\partial\Omega$ and satisfy the following assumption
Assumption 1.1.
  1. 1.
    $\partial\Omega_{1}\cup\partial\Omega_{2}=\partial\Omega$.
  2. 2.
    $\partial\Omega_{1}$ and $\partial\Omega_{2}$ are of positive measure.
  3. 3.
    $\partial\Omega_{1}$ and $\partial\Omega_{2}$ are unions of connected components of $\partial\Omega$.
  4. 4.
    $\partial\Omega_{1}$ is relative open to $\partial\Omega$.
  5. 5.
    $\overline{\partial\Omega_{1}}\cap\overline{\partial\Omega_{2}}$ consists of finitely many $C^{2}$ curves in $\mathbb{R}^{3}$.
Here $m_{T}=\frac{1}{2\pi T^{2}}e^{-\frac{|v|^{2}}{2T}}$ and $m_{T(x)}=\frac{1}{2\pi T^{2}(x)}e^{-\frac{|v|^{2}}{2T(x)}}$. Also, we denote $m=m_{1}$. Let $\partial E$ is a subset of $\partial\Omega$. We define $\partial E^{-},\partial E^{+}$ and $\partial E^{0}$ as follows.
$\partial E^{-}$ $=\{(x,v)\in\partial E\times\mathbb{R}^{3}:\,n(x)\cdot v<0\},$
$\partial E^{+}$ $=\{(x,v)\in\partial E\times\mathbb{R}^{3}:\,n(x)\cdot v>0\},$
$\partial E^{0}$ $=\{(x,v)\in\partial E\times\mathbb{R}^{3}:\,n(x)\cdot v=0\},$
where $n(x)$ denotes the outward unit normal vector as $x\in\partial\Omega$. In this article, we always assume that $T(x)$ is positive and belongs to $L^{\infty}(\partial\Omega)$. Furthermore, we assume that $\mathbf{B}(|v-u|,\omega)$ satisfies
$\mathbf{B}(|v-u|,\omega)=\mathfrak{B}|v-u|^{\gamma}\left(\frac{u-v}{|u-v|}\cdot\omega\right),\quad-3<\gamma\leq 1,$ (1.6)
for some constant $\mathfrak{B}>0$. When we discuss the thermal transpiration, we define two parameters as
$A$ $\coloneq\inf\{x_{3}\in\mathbb{R}^{3}|(x_{1},x_{2},x_{3})\in\Omega\},$
$B$ $\coloneq\sup\{x_{3}\in\mathbb{R}^{3}|(x_{1},x_{2},x_{3})\in\Omega\}.$
Let $A<a<b<B$, and define
$\partial\Omega_{1}=\Big(\partial\Omega\cap\{z<a\}\Big)\cup\Big(\partial\Omega\cap\{z>b\}\Big).$ (1.7)
Finally, we set the distribution of gas $F(x,v)$ in $\Omega$ which satisfies the following boundary value problem:
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (1.8)
Furthermore, we impose that $T(x_{1},x_{2},x_{3})=1$ when $x_{3}\leq a$, and that $T(x_{1},x_{2},x_{3})=T_{2}$ when $x_{3}\geq b$. We write the solution as a perturbation near the standard Maxwellian $M(v)=\frac{1}{\sqrt{2\pi}}m(v)$. Let $F=M+m^{1/2}f$ and $F_{in}=M+m^{1/2}f_{in}$. Then, the function $f$ solves the following boundary value problem:
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=-\frac{1}{\kappa}L(f)+\frac{1}{\kappa}\Gamma(f,f),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f(x,v)&=f_{in}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f(x,v)&=m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (1.9)
where
$L(f)$ $=m^{-1/2}\Big(Q(M,m^{1/2}f)+Q(m^{1/2}f,M)\Big),$
$\Gamma$ $=m^{-1/2}Q(m^{1/2}f,m^{1/2}f),$
$P(f)(x,v)$ $=m^{1/2}\int_{u\cdot n(x,v)>0}f(x,u)m^{1/2}u\cdot n(x)du.$
The operator $L$ has two components: the multiplicative operator $\nu$ and integral operator $K$, that is , $L(f)=\nu(v)f-K(f)$. Their definition can be found in Chapter 2 or [17]. In what follows, we discuss the well-posedness of the boundary value problem (1.9). The integral form of (1.9) is
$f=\frac{1}{\kappa}S^{\kappa}_{\Omega}Kf+J^{\kappa}\left(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+\left(m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\right)\textbf{1}_{\partial\Omega^{-}_{2}}\right)+\frac{1}{\kappa}S^{\kappa}_{\Omega}\Gamma(f,f).$ (1.10)
Here, we define
$S_{\Omega}^{\kappa}f(x,v)$ $\coloneq\int_{0}^{\tau_{-}(x,v)}e^{-\frac{1}{\kappa}\nu(v)s}f(x-sv,v)ds,$ (1.11)
$J^{\kappa}f$ $\coloneq e^{-\frac{1}{\kappa}\nu(v)\tau_{-}(x,v)}f(q(x,v),v),$ (1.12)
where
$\tau_{-}(x,v)=\sup\{t\,|\,x-tv\in\overline{\Omega}\}\quad\text{and}\quad q(x,v)=x-\tau_{-}(x,v)v.$ (1.13)
For convenience, we define
$\langle f,g\rangle$ $=\int_{\Omega\times\mathbb{R}^{3}}fgdxdv,$ (1.14)
$\langle f,g\rangle_{\partial^{-}_{2}}$ $=\int_{\partial\Omega^{-}_{2}}fg|v\cdot n(x)|dxdv.$ (1.15)
To state our main theorem, we introduce some function spaces. For $\alpha\geq 0$, $\beta\in\mathbb{R}$ and $p\geq 1$, we say $f\in L^{p}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})$ if
$\begin{aligned} \lVert f\rVert_{L^{p}_{\alpha,\beta}}=\lVert f\rVert_{L^{p}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}&\coloneq\left(\int_{\Omega}\int_{\mathbb{R}^{3}}\left|f(x,v)e^{\alpha|v|^{2}}(1+|v|)^{\beta}\right|^{p}dvdx\right)^{1/p}<\infty\quad&&\text{for }p<\infty,\\ \lVert f\rVert_{L^{\infty}_{\alpha,\beta}}=\lVert f\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}&\coloneq\underset{{(x,v)\in\Omega\times\mathbb{R}}}{\text{ess sup}}\left|f(x,v)e^{\alpha|v|^{2}}(1+|v|)^{\beta}\right|<\infty&&\text{for }p=\infty.\end{aligned}$
Moreover, we say $f\in L^{p}_{\alpha,\beta}(\partial E^{-})$ for some $\partial E\subset\partial\Omega$, if
$\begin{aligned} \lVert f\rVert_{L^{p}_{\alpha,\beta}(\partial E^{-})}&\coloneq\left(\int_{\partial E^{-}}\left|f(x,v)e^{\alpha|v|^{2}}(1+|v|)^{\beta}\right|^{p}|v\cdot n(x)|dSdv\right)^{1/p}<\infty\quad&&\text{for }p<\infty,\\ \lVert f\rVert_{L^{\infty}_{\alpha,\beta}(\partial E^{-})}&\coloneq\underset{{(x,v)\in\partial E^{-}}}{\text{ess sup}}\left|f(x,v)e^{\alpha|v|^{2}}(1+|v|)^{\beta}\right|<\infty&&\text{for }p=\infty,\end{aligned}$
where $dS$ denotes the surface measure on $\partial\Omega$. For convenience, we write $L^{p}(A)=L^{p}_{0,0}(A)$ when $A=\Omega\times\mathbb{R}^{3}\,\text{or }\partial E^{-}\subseteq\partial\Omega\times\mathbb{R}^{3}$. Throughout this article, we denote $f(x)=\mathcal{O}(g(x))$ when $\lim_{x\to\infty}\frac{f(x)}{g(x)}$ is a constant.
We now introduce the concept of an unsealed set $\Omega$ in $\mathbb{R}^{3}$. The term unseal describes the visibility of $\partial\Omega_{1}$ from points on $\partial\Omega$. If every point $x$ can directly see the $\partial\Omega_{1}$, then we call $\Omega$ a completely unsealed set. If, for every $x\in\partial\Omega$, the $\partial\Omega_{1}$ becomes visible after tracing back $m$ times for some fixed integer $m$, then we call $\Omega$ an unsealed set. The precise definition is in Chapter 4.
Theorem 1.1.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Suppose that $\Omega$ is unsealed with respect to $\partial\Omega_{1}$ and satisfies the assumption 1.1. Assume that $3<\gamma\leq 1$ and that (1.6) holds. Then, for any $0<\alpha<1/4$ and $\kappa_{0}\in(0,\infty)$, there exists $\delta>0$ which depends on $\alpha$, $\Omega$, $\partial\Omega_{1}$, and $\kappa_{0}$ such that if
$\lVert m^{-1/2}\left(F_{in}-M\right)\rVert_{L^{\infty}_{a,0}(\partial\Omega^{-}_{1})}\leq\delta$ (1.16)
and
$|T(x)-1|\leq\delta,\quad\forall x\in\Omega_{2}$ (1.17)
then the equation 1.5 has a unique mild solution for $\kappa\geq\kappa_{0}$. Furthermore,
$\lVert f\rVert_{L^{\infty}_{\alpha,0}}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}\Big(\left\lVert m^{-1/2}\left(F_{in}-M\right)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\left\lVert m^{-1/2}(m_{T(x)}-m)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big).$ (1.18)
Here, $f=m^{-1/2}\left(F-M\right)$.
For $x=(x_{1},x_{2},z)\in\Omega$, the local flux $u(x)$ and the total flux $U(z)$ in the $x_{3}$-direction are defined as follows:
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{(x_{1},x_{2},z)\in\Omega\}}u(x)dS.$ (1.19)
By conservation law, the total flux is a constant with respect to $z$ when $z\in[a,b]$.
Theorem 1.2.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose for every $\kappa>\kappa_{0}$, there exists a constant $\mathcal{C}$ such that (1.8) admits a mild solution with the following estimate:
$\lVert F\rVert_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}\leq\mathcal{C},$ (1.20)
where $\beta>\max\{\gamma,0\}+5$. Then, there exists $\kappa_{1}>\kappa_{0}$ such that the total flux $U(x_{3})$ of $F$ is positive whenever $\kappa\geq\kappa_{1}$. Moreover,
$U(x_{3})\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (1.21)
In Knudsen’s experiment ([25]), he connects two chambers with different temperature $T_{1},T_{2}$ with a narrow tube. He measures the pressures in two tanks $P_{1},P_{2}$ at equilibrium when there is no flux in the tube. He observes in the limit of infinitely small tubing, $\frac{P_{1}}{\sqrt{T_{1}}}=\frac{P_{2}}{\sqrt{T_{2}}}$. For the free transport equation case, the relation is exact ([36]). In our case, we obtain the same result by considering the following boundary value problem:
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (1.22)
Theorem 1.3.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary such that $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex. Suppose that for every $\kappa>\kappa_{0}$, equation (1.8) admits a mild solution bounded in $L^{\infty}_{0,\beta}$ by a constant independent of $\kappa$, where $\beta>\max\{\gamma,0\}+5$. Then, the total flux $U(x_{3})$ of $F$ satisfies the following estimate:
$|U(x_{3})|=\mathcal{O}(\frac{1}{\kappa}).$ (1.23)
We can extend Theorem 1.3 to the case when the incoming data is not localized at two places, i.e.,
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=\frac{1}{\kappa}Q(F,F),&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{i}}e^{-\frac{v^{2}}{2}},&&\text{on}\quad\partial B^{-}_{i},i=1,\cdots,n,\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\cup_{i=1}^{n}\partial B^{-}_{i}\right),\end{aligned}\right$. (1.24)
where $\partial B_{i}$ is an open subset of $\partial\Omega$ and we assume that $T(x)=T_{i}$ if $x_{i}\in\partial B_{i}$.
Remark 1.1.
Note that the condition $\Omega\cap\{a\leq x_{3}\leq b\}$ is convex ensures that $\Omega$ is an unsealed set with respect to
$\Big(\partial\Omega\cap\{z<a\}\Big)\cup\Big(\partial\Omega\cap\{z>b\}\Big).$
The proof of Theorem 1.2 and Theorem 1.3 is valid for a domain in which $\Omega\cap\{a\leq x_{3}\leq b\}$ is not convex but is unsealed with respect to $\Big(\partial\Omega\cap\{z<a\}\Big)\cup\Big(\partial\Omega\cap\{z>b\}\Big)$.
Remark 1.2.
Note that when $\Omega$ is a bounded convex domain with $C^{2}$ boundary. The Theorem 1.1 ensures the condition of Theorem 1.2.
1.2. Organization of the article
We briefly outline the proof in this article. We use the traditional $L^{2}-L^{\infty}$ estimate together with the bootstrap method to construct a solution of (1.5).
In Chapter 2, we review some elementary properties of the linear operator $L$. In Chapter 3, inspired by [13], [24], we introduce a simpler and more efficient approach to obtaining the $L^{2}$ estimate. This approach is based on the operator $S_{\Omega}^{\kappa}K$, which is compact in $L^{2}_{0,\gamma/2}$.
In Chapter 4, we use the unsealed condition to obtain the $L^{\infty}$ estimate. This concept provides several advantages, such as reducing the weight requirement in $L^{\infty}$. Moreover, this is a key tool in proving Theorem 1.2 and Theorem 1.3.
In Chapter 5, we use original Boltzmann equation to address the problem by modeling the distribution as in (1.8), with incoming data at $\Big(\partial\Omega\cap\{z<a\}\Big)$ and $\Big(\partial\Omega\cap\{z>b\}\Big)$. We cannot apply the powerful technique developed in [6]. However, we observe that the crucial mechanism for the appearance of flux lies in the different between the incoming data and the solution of the free transport model, which satisfies (1.6). By combining this observation with the concept of the unsealed set, we complete the proof of the main theorems in this article.
Chapter 2. Preliminaries
In this chapter, we review some properties of the linear operator $L$ and its components: the multiplication operator $\nu$ and the integral operator,$K$. We also present several propositions about the operator $S^{\kappa}_{\Omega}$ and $J^{\kappa}$, which are useful in Chapter 3 and Chapter 4. In addition, we recall some basic properties of $Q_{+}$ and $Q_{-}$.
We emphasize one important point: Kawagoe Daisuke [24] and others normalize the Maxwellian as $m=e^{-|v|^{2}}$. Thus the exponents differ by a factor of 2 and the constants are slightly different.
2.1. Estimates for $\nu$ and $K$
Recall that $L(f)=\nu(v)f-K(f)$ when the assumption (1.6) holds, and that $\nu(v)$ and $K(f)$ have explicit forms as follows
$\nu(v)$ $=\iint_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-u|,\sigma)m(u)dud\sigma,$
$K(f)(x,v)$ $=\int_{\mathbb{R}^{3}}k(v,u)f(x,u)du=\int_{\mathbb{R}^{3}}\big(k_{1}(v,u)-k_{2}(v,u)\big)f(x,u)du.$
The explicit formulas for $k_{1}$ and $k_{2}$ are:
$k_{1}(v,u)$ $=B_{0}\pi|v-u|^{\gamma}e^{-\frac{1}{4}(|v|^{2}+|u|^{2})},$
$k_{2}(v,u)$ $=\frac{2B_{0}}{\pi^{3/2}}\frac{1}{|v-u|}e^{-\frac{1}{8}|v-u|^{2}-|V_{1}(v,u)|^{2}}\times\int_{W_{v-u}}e^{-|w+V_{2}(v,u)|^{2}}\left(|v-u|^{2}+|w|^{2}\right)^{\frac{\gamma}{2}}dw,$
where
$V_{1}(v,u)$ $\coloneq\frac{1}{2\sqrt{2}}\frac{|v|^{2}-|u|^{2}}{|v-u|^{2}}(v-u),$
$V_{2}(v,u)$ $\coloneq\frac{(v-u)\times(v\times u)}{\sqrt{2}|v-u|^{2}},$
$W_{v-u}$ $\coloneq\{w\in\mathbb{R}^{3}|w\perp(v-u)\}.$
Notice that we have the following identity
$|V_{1}(v,u)|^{2}+|V_{2}(v,u)|^{2}$ $=\frac{1}{8}|v+u|^{2},$
$V_{2}(v,u)\cdot(v-u)$ $=0,$
$\frac{1}{8}|v-u|^{2}+|V_{1}(v,u)|^{2}+|V_{2}(v,u)|^{2}$ $=\frac{1}{4}(|v|^{2}+|u|^{2}).$
For their derivation, see [12]. We first give some estimates for the collision frequency $\nu(v)$.
Proposition 2.1.
Suppose that (1.6) holds and $3<\gamma\leq 1$. For the collision frequency $\nu(v)$, there exist positive constant $\nu_{0}$ and $\nu_{1}$ such that
$\nu_{0}(1+|v|)^{\gamma/2}\leq\nu(v)\leq\nu_{1}(1+|v|)^{\gamma/2}$
for all $v\in\mathbb{R}^{3}$.
For a proof of Proposition 2.1, see [12], [17].
Proposition 2.2.
Let $3<\gamma\leq 1$ and let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$. Suppose (1.6) holds. Then, we have
$\int_{0}^{\frac{L}{|v|}}e^{-\frac{1}{\kappa}\nu(|v|)t}dt\leq C_{\gamma,L}\frac{\kappa}{1+|v|}$
for any fix $L>0$ and $(x,v)\in\Omega\times\mathbb{R}^{3}$.
The proof of Proposition 2.2 can be found in [44]. We next introduce some estimate for integral operator $K(f)$.
Proposition 2.3.
Let $3<\gamma\leq 1$ and $0<\delta<1$. Suppose that (1.6) holds. Then, we have
$|k(v,u)|\leq C_{\gamma,\delta}\frac{w_{\gamma}(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}E_{\delta}(v,u)$ (2.1)
for all $v,u\in\mathbb{R}^{3}$, where
$E_{\delta}(v,u)\coloneq e^{-\frac{1-\delta}{8}\left(|v-u|^{2}+4|V_{1}(v,u)|^{2}\right)}=e^{-\frac{1-\delta}{8}\left(|v-u|^{2}+\left(\frac{|v|^{2}-|u|^{2}}{|v-u|}\right)^{2}\right)}$
and
$w_{\gamma}(|v-u|)\coloneq\left\{\begin{aligned} &\frac{1}{|v-u|},&&-1<\gamma\leq 1,\\ &\frac{|\ln|v-u||+1}{|v-u|},&&\gamma=-1,\\ &\frac{1}{|v-u|^{|\gamma|}},&&-3<\gamma<-1.\end{aligned}\right$. (2.2)
The detail of proof can be found in [12] and the appendix of [24].
In the chapter 4, we estimate $L^{\infty}_{\alpha,0}$ norm of the solution of Boltzmann equation. Hence, we need to discuss
$K_{\alpha}(f)\coloneq w(v)K(w^{-1}f)=\int_{\mathbb{R}^{3}}w(v)k(v,u)w^{-1}(u)f(x,u)du\coloneq\int_{\mathbb{R}^{3}}k_{\alpha}(v,u)f(x,u)du$
for $w(v)=e^{\alpha|v|^{2}}$, $\alpha>0$. It is known in [8] that
$E_{\delta}(v,u)$ $=e^{a|v|^{2}}e^{-\alpha_{1,a,\delta}|v-u|^{2}}e^{-(1-\delta)\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2,a,\delta}|v-u|\right)^{2}}e^{-a|u|^{2}}$
$=e^{-a|v|^{2}}e^{-\alpha_{1,a,\delta}|v-u|^{2}}e^{-(1-\delta)\left(\frac{(v-u)\cdot v}{|v-u|}+\alpha_{2,a,\delta}|v-u|\right)^{2}}e^{a|u|^{2}}$
for 1/4<a<1/4-1/4<a<1/4, $\delta>0$ and $v,u\in\mathbb{R}^{3}$, where
$\alpha_{1,a,\delta}$ $\coloneq\frac{(1-\delta)^{2}-4a^{2}}{8(1-\delta)},$
$\alpha_{2,a,\delta}$ $\coloneq\frac{1-\delta-2a}{2\sqrt{2}(1-\delta)}.$
Plugging these identities into that estimate 2.1 with $a=\alpha$ and $\delta=1/2-\alpha$, we obtain
$|k_{\alpha}(v,u)|\leq C_{\gamma,\alpha}\frac{w(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$ (2.3)
for some $\alpha_{1},\alpha_{2},\alpha_{3}>0$ depending only on $\alpha$. Based on the above calculations, we obtain the following estimates, which are related to the kernel of the operator $K_{\alpha}$.
Proposition 2.4.
[4] Let $\alpha_{1},\alpha_{3}>0$ and $\alpha_{2}\in\mathbb{R}^{3}$. Then, we have
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq\frac{C_{\alpha,\mu}}{1+|v|}$ (2.4)
for any $\mu<3$ and $v\in\mathbb{R}^{3}$.
Proof.
Let w=vuw=v-u. By using the sphere coordinate and setting $\omega=|v|\cos\theta-\alpha_{2}r$, we have
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}dv$ $=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}\left(|v|\cos\theta-\alpha_{2}r\right)^{2}}\sin\theta d\theta d\phi dr$
$=\frac{2\pi}{|v|}\int_{0}^{\infty}\int_{-|v|-\alpha_{2}r}^{|v|-\alpha_{2}r}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}w^{2}}d\omega\phi dr$
$\leq C_{\alpha_{1},\alpha_{3}}\frac{1}{|v|}.$
Moreover, we can deduce that
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq C.$
Here, the constant $C$ is independent to $v$. combining these results, the proof is complete. ∎
Proposition 2.5.
[24] Let $0\leq\alpha<1/4,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $1\leq q<\min\{3,\frac{3}{|\gamma|}\}$. Suppose that (1.6) holds. Then we have
$\displaystyle\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du\leq C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$ (2.5)
for all $v\in\mathbb{R}^{3}.$
Remark 2.1.
We set $\min\{3,\frac{3}{|\gamma|}\}=3$ when $\gamma=0$
The proof can be found in [24]. For convenience, we present the proof here.
Proof.
Case 1 $\beta\geq 0$:
Notice that $(1+|u|)^{\beta}\leq C_{\beta}\left((1+|v|)^{\beta}+|v-u|^{\beta}\right)$ for all $v,u\in\mathbb{R}^{3}$. By 2.3, we have
$|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}$ $\leq C(1+|v|)^{\beta+q(\gamma-1)}w_{\gamma}(|v-u|)^{q}e^{-q\alpha_{1}|v-u|}e^{-q\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$
$\quad+C(1+|v|)^{q(\gamma-1)}w_{\gamma}(|v-u|)^{q}|v-u|^{\beta}e^{-q\alpha_{1}|v-u|}e^{-q\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}.$
Observe that there exists $\mu_{\gamma,q}<3$ such that $w_{\gamma}(|v-u|)^{q}\leq|v-u|^{-\mu_{\gamma,q}}$. The constant $\mu_{\gamma,q}$ only depend on $\gamma$ and $q$. Also, we note
$w_{-1}(|v-u|)e^{-\alpha_{1}|v-u|^{2}}\leq C_{\epsilon}|v-u|^{-(1+\epsilon)}$ (2.6)
for some small positive constant $\epsilon$. Thus, we can treat the case $\gamma=-1$ in the same way as for $3<\gamma<-1$. In any case, by (2.4) we have
$\displaystyle\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du$ $\leq C_{\alpha,\gamma}(1+|v|)^{q(\gamma-1)-1}+C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$
$\leq C_{\alpha,\gamma}(1+|v|)^{q(\gamma-1)-1}.$
Case 2 $\beta<0$:
Observe that
$(1+|u|)^{\beta}$ $=(1+|v|)^{\beta}\frac{(1+|v|)^{|\beta|}}{(1+|u|)^{|\beta|}}$
$\leq(1+|v|)^{\beta}\frac{(1+|u|)^{|\beta|}+|v-u|^{|\beta|}}{(1+|u|)^{|\beta|}}$
$\leq(1+|v|)^{\beta}\left(1+|v-u|^{|\beta|}\right).$
We apply the above argument to obtain the same estimate. Thus, the proof is complete. ∎
2.2. Proposition of $L$
Next, we review some basic properties of $L$, including the spectral gap and the projection of $L$. Recall that the normal orthogonal basis of the null space Ker $L$ is given by
$\Phi_{0}$ $=(2\pi)^{-\tfrac{3}{4}}\,m^{\tfrac{1}{2}}(v),$
$\Phi_{i}$ $=(2\pi)^{-\tfrac{3}{4}}v_{i}m^{\tfrac{1}{2}}(v),\quad i=1,2,3,$
$\Phi_{4}$ $=\frac{(2\pi)^{-\tfrac{3}{4}}}{\sqrt{6}}\,(|v|^{2}-3)m^{\tfrac{1}{2}}(v).$
For each $f=f(v)\in L^{2}(\mathbb{R}^{3}_{v})$, we denote the macroscopic part $P_{L}(f)$ as the projection of $f$ onto Ker $L$, that is,
$P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i},$
and further denote $(I-P_{L})(f)=f-P_{L}(f)$ to be the microscopic part of $f$. It is well-known that there exists a constant $C_{L}$ such that
$\langle L(f),f\rangle_{v}\geq C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}(dv)}.$ (2.7)
Here $L^{2}_{0,\gamma/2}(dv)$ denotes integration only with respect to the variable $v$. The proof can be found in [17], [20] and [22]. The most important estimate concerning $P_{L}$ and $(I-P_{L})$ is the following.
Proposition 2.6.
Let $\Omega$ be a bounded connected domain in $\mathbb{R}^{3}$ with $C^{2}$ boundary and $3<\gamma\leq 1$. Suppose that $f\in L^{2}_{0,\gamma/2}$ is a solution to
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}\end{aligned}\right$. (2.8)
in distribution sense. Then,
$\lVert P_{L}(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert(I-P)f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}\end{pmatrix}.$ (2.9)
Proof.
Recall $P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}$. Thus, it suffices to compute $\langle f,\Phi_{i}\rangle$, for $i=0,\cdots,4$. The estimate of $\langle f,\Phi_{i}\rangle$ are the same as those given in lemma 3.3 in [14] for $i=1,\cdots,4$. The standard elliptic estimates appearing in [14] can also be found in the chapter 2 of [18] or chapter 6 of [15]. For $i=0$, we choose a test function
$\psi=\sum_{j=1}^{3}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{j}\partial_{j}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0},$ (2.10)
where $\phi_{\langle f,\Phi_{0}\rangle_{v}}$ is the solution of following elliptic problem
$\left\{\begin{aligned} -\Delta\phi_{\langle f,\Phi_{0}\rangle_{v}}&=\langle f,\Phi_{0}\rangle_{v},&&\text{in }\Omega,\\ \hat{\gamma}(x)\phi_{\langle f,\Phi_{0}\rangle_{v}}+\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (2.11)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ ( but $\gamma>0$ on a positive measure subset of $\partial\Omega$ ). Notice that we have following estimate
$\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(\Omega)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(\Omega)}.$ (2.12)
This estimate can be found in [16] and [34]. We choose $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ such that
$\int_{\mathbb{R}^{3}}\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)\left(\frac{|v|^{2}}{2}-\frac{3}{2}\right)(v_{i})^{2}\,\mu(v)\,dv=0.$
Then, following the same discussion as in [14], we obtain
$\lVert\langle f,\Phi_{0}\rangle_{v}\rVert^{2}_{L^{2}(\Omega)}\leq C\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}(\Omega\times\mathbb{R}^{3})}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}+\lVert(I-P)(f)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\end{pmatrix}.$ (2.13)
Combining all these estimates, the proof is completed. ∎
2.3. Proposition of $S_{\Omega}$ and $J$
Proposition 2.7.
Let $\Omega$ be a bounded convex domain with $C^{1}$ boundary, $\alpha\geq 0,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $2\leq p\leq\infty$. Suppose that (1.6) holds. Then, the operator norm of $S^{\kappa}_{\Omega}:L^{p}_{\alpha,\beta}\to L^{p}_{\alpha,\beta+1}$ is $\mathcal{O}(\kappa)$.
Proposition 2.8.
[…]
$\lVert J^{\kappa}\large(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}})\rVert_{L^{2}_{0,\gamma/2}}\leq C\kappa\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}\Big).$ (2.14)
Proof.
By Lemma 2.1 ([7]), we have
$\lVert J^{\kappa}\large(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}})\rVert^{2}_{L^{2}_{0,\gamma/2}}$
$=$ $\int_{\Omega}\int_{\mathbb{R}^{3}}e^{-2\frac{1}{\kappa}\nu(|v|)\tau_{-}(x,v)}\Big[\big(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}}\big)(q(x,v),v)\Big]^{2}(1+|v|)^{\gamma}dvdx$
$=$ $\int_{\Omega^{-}}\int_{0}^{\tau_{-}(x,-v)}e^{-2\frac{1}{\kappa}\nu(|v|)t}\Big[\big(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}}\big)(z,v)\Big]^{2}(1+|v|)^{\gamma}|v\cdot n(x)|dtdSdv$
$=$ $\kappa\int_{\Omega^{-}}\Big[\big(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}}\big)(z,v)\Big]^{2}\frac{(1+|v|)^{\gamma}}{\nu(|v|)}|v\cdot n(x)|dSdv$
$\leq$ $C\kappa\int_{\Omega^{-}}\big(f^{2}_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+P^{2}(f)\textbf{1}_{\partial\Omega^{-}_{2}}+\phi^{2}_{2}\textbf{1}_{\partial\Omega^{-}_{2}}\big)(z,v)|v\cdot n(x)|dSdv$
[…]
This finish the proof. ∎
2.4. Estimate of $Q_{+}$ and $Q_{-}$
Lemma 2.1.
For any $\beta\geq 1$,
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{-\beta}d\sigma\leq C_{\beta}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-(\beta+1)},$ (2.15)
where $v^{\prime}=\frac{v+u}{2}+\frac{|v-u|}{2}\sigma$ and $u^{\prime}=\frac{v+u}{2}-\frac{|v-u|}{2}\sigma$.
Proof.
For convenience, let $A=1+\left|\frac{v+u}{2}\right|^{2}+\left|\frac{v-u}{2}\right|^{2}$ and $B=2\left|\frac{v+u}{2}\right|\left|\frac{v-u}{2}\right|$. When $\beta=1$,
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{\beta}d\sigma=\int_{0}^{2\pi}\int_{0}^{\pi}\frac{\sin\theta}{\sqrt{A^{2}-B^{2}\cos^{2}\theta}}d\theta\phi.$
Here we use sphere coordinate with $e_{1}=\frac{v+u}{|v+u|}$. Let $z=\cos\theta$ and $y=\sin^{-1}\left(Bz/A\right)$. We obtain
$\int_{0}^{2\pi}\int_{0}^{\pi}\frac{\sin\theta}{\sqrt{A^{2}-B^{2}\cos^{2}\theta}}d\theta\phi$ $=4\pi\int_{0}^{1}\frac{1}{\sqrt{A^{2}-B^{2}z^{2}}}dz$
$=4\pi\int_{0}^{\sin^{-1}(B/A)}\frac{\cos(y)}{B|\cos(y)|}dy$
$\leq 8\pi A^{-1}$
$\leq\frac{16\pi}{1+|v|^{2}+|u|^{2}}.$
The penultimate inequality follows from $\sin^{-1}(x)\leq 2x$ for any $0\leq x\leq 1$. For $\beta>1$, we first consider the case of $\beta\in\mathbb{N}$. We use induction to establish the result. Since $\beta=1$ holds by the above calculation. assume that $\beta=n$ holds. When $\beta=n+1$, we have
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-(n+1)}\langle u^{\prime}\rangle^{-(n+1)}d\sigma$ $=4\pi\int_{0}^{B/A}BA^{-n}\frac{1}{\left(\sqrt{1-y^{2}}\right)^{n+1}}dy$
$\leq\frac{4\pi}{A\sqrt{1-(B/A)^{2}}}\int_{0}^{B/A}BA^{1-n}\frac{1}{\left(\sqrt{1-y^{2}}\right)^{n}}dy$
$\leq\frac{C}{\sqrt{A^{2}-B^{2}}}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{n+1}$
$\leq\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{n+2}.$
The second-to-last line follows from the induction hypothesis, so (2.15) holds when $\beta\in\mathbb{N}$. For $\beta\notin\mathbb{N}$, Suppose that $\beta=n+d$ for some $n\in\mathbb{N}$ and $d\in(0,1)$. Then, we have
$4\pi\int_{0}^{B/A}BA^{1-\beta}\frac{1}{\left(\sqrt{1-y^{2}}\right)^{\beta}}dy$ $\leq\frac{4\pi}{\left(\sqrt{A^{2}-B^{2}}\right)^{d}}\int_{0}^{B/A}BA^{1-n}\frac{1}{\left(\sqrt{1-y^{2}}\right)^{n}}dy$
$\leq\frac{C}{\left(\sqrt{A^{2}-B^{2}}\right)^{d}}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-n-1}$
$=C\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-(\beta+1)}.$
Now, we provide an estimate of $Q_{+},Q_{-}$, which will be used in Chapter 5.
Proposition 2.9.
Suppose that (1.6) holds. Then,
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (2.16)
for any $\beta>5+\max\{0,\gamma\}$ and
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{-}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (2.17)
for any $\beta>3+\max\{0,\gamma\}$.
Proof.
The proof of 2.17 follows directly from
$\lVert S^{\kappa}(f)\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}\leq\frac{C}{|v|}\lVert f\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}.$ (2.18)
So it is omitted here for simplicity of presentation. The proof of (2.16) proceeds as follows:
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|$ $\leq C\left|\int_{\mathbb{R}^{3}}\lVert Q_{+}[f,f]\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}dv\right|$
$\leq\int_{\mathbb{R}^{3}}\int_{\mathbb{R}^{3}}\int_{\mathbb{S}^{2}}|v-u|^{\gamma}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{-\beta}d\sigma dudv\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$
$\leq C\int_{\mathbb{R}^{3}}\int_{\mathbb{R}^{3}}|v-u|^{\gamma}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-(\beta+1)}dudv\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}.$
The second-to-last inequality follows from Lemma (2.1). ∎
Chapter 3. $L^{2}$ estimate
Before we begin the $L^{2}$ estimate, we recall a lemma concerning boundary data. This lemma ensures that the trace of $f$ is well-defined locally for a certain class of functions. The proof can be found in [14].
Lemma 3.1 (Ukai Trace Theorem).
Define
$\gamma^{\varepsilon}\equiv\{(x,v)\in\gamma:|n(x)\cdot v|\geq\varepsilon,\ |v|\leq\tfrac{1}{\varepsilon}\}.$ (3.1)
Then
$\|f1_{\gamma^{\varepsilon}}\|_{L^{1}(\partial\Omega\times\mathbb{R}^{3})}\;\leq C_{\varepsilon,\Omega}\Big(\|f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}+\|v\cdot\nabla_{x}f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}\Big).$ (3.2)
Also, we recall Green’s identity. The proof can be found in [5].
Lemma 3.2.
Assume that $f(x,v),\,g(x,v)\in L^{2}(\Omega\times\mathbb{R}^{3})$, $v\cdot\nabla_{x}f,\;v\cdot\nabla_{x}g\in L^{2}(\Omega\times\mathbb{R}^{3})$, and $f_{\gamma},\;g_{\gamma}\in L^{2}(\partial\Omega\times\mathbb{R}^{3})$. Then
$\iint_{\Omega\times\mathbb{R}^{3}}\bigl\{(v\cdot\nabla_{x}f)\,g+(v\cdot\nabla_{x}g)\,f\bigr\}\,dv\,dx=\int_{\gamma}fg\,d\gamma.$ (2.3)
We now state the main theorem of this chapter. The statement and strategy are similar to [14] and [13]. However, we use the Lemma 1.2 ([24]) to simplify the process. For convenience, we present the detail here.
Theorem 3.1.
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Assume that $3<\gamma\leq 1$ and that (1.6) holds. Then, for any $\phi_{1}\in L^{2}_{0,-\gamma/2}(\Omega\times\mathbb{R}^{3})$, $\phi_{2}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{1})$, and $f_{in}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{2})$. There exists a unique solution $f\in L^{2}_{0,\gamma/2}$ to the following boundary value problem.
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (3.3)
Moreover,
[…]
Proof.
Consider the following iteration with $f_{1}=0$,
$\left\{\begin{aligned} \hat{\alpha}\nu(v)f_{k+1}+v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})&=\phi_{1},&&\quad\text{in}\quad\in\Omega\times\mathbb{R}^{3},\\ f_{k+1}&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}&=\eta P(f_{k})+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (3.5)
where $0<\hat{\alpha},\eta<1$. By Green formula, we have
$\hat{\alpha}\lVert f_{k+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f_{k+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f_{k+1}),f_{k+1}\rangle$ (3.6)
$=\langle\phi_{1},f_{k+1}\rangle+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f_{k})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
In particular, we have
[…]
Using Proposition 2.8 , we deduce that
$J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})\in L^{2}_{0,\gamma/2}\qquad\text{if }\,f_{k}\in L^{2}_{\Omega^{+}_{2}}.$ (3.8)
We can then rewrite (3.5) in integral form as:
$(I-\frac{1}{\kappa}S^{\kappa}_{\Omega}K)f_{k+1}=J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})+S_{\Omega}\phi_{1}.$ (3.9)
Applying Lemma 1.2 ([24]), we establish the existence of $f_{k+1}$. By induction, we can then prove that $f_{k+1}$ exists and belongs to $L^{2}_{0,\gamma/2}$ for any $k\in\mathbb{N}$. Applying Yong’s inequality, we have
$\lVert\eta P(f_{k})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\eta^{2}\lVert P(f_{k})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+2\eta\langle P(f_{k+1}),\phi_{2}\rangle_{\partial^{-}_{2}}$ (3.10)
$\leq\eta\lVert P(f_{k})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\frac{1}{1-\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
$\leq\eta\lVert f_{k}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\frac{1}{1-\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
Then, we combine the (3.6) and (3.10) to obtain
[…]
$\quad+C_{\alpha}\left(\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\right)+C_{\hat{\alpha},\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
By same process, we have
$\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $+(2\hat{\alpha})^{-1}\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ (3.12)
[…]
This implies that $\{f_{k}\}$ is a Cauchy sequence. For any fixed $\eta$ and $\hat{\alpha}$, let $\lim_{k\to\infty}f_{k}\coloneqq f^{\hat{\alpha},\eta}$. By applying Green’s formula once more, we have
$\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f^{\hat{\alpha},\eta}),f^{\hat{\alpha},\eta}\rangle$ (3.13)
$=\langle\phi_{1},f^{\hat{\alpha},\eta}\rangle+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\hat{\alpha},\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
Applying (2.9) we have
$\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\\ &+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.\end{pmatrix}$. (3.14)
Multiplying it with a constant $a>0$ and combining it with equation (3.13), we derive that
$\hat{\alpha}(1-aC)\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+a\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+C_{L}\frac{1}{\kappa}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$
[…]
$+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\hat{\alpha},\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\langle\phi_{1},f^{\hat{\alpha},\eta}\rangle.$
Finally, by choosing $a$ sufficiently small, we conclude that
$\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}$ $+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (3.15)
[…]
Hence, $f^{\hat{\alpha},\eta}$ is a bounded sequence in $L^{2}_{0,\gamma/2}$ for any fixed $\eta$. Observe that $f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{1}}$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)+\frac{1}{\kappa}L\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)&=\hat{\alpha}_{1}\nu f^{\eta,\hat{\alpha}_{1}}-\hat{\alpha}_{2}\nu f^{\eta,\hat{\alpha}_{2}},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}&=\eta P\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$.
for any $0<\hat{\alpha}_{1},\hat{\alpha}_{2}<1$. After taking $\hat{\alpha}_{1}\nu f^{\eta,\hat{\alpha}_{1}}-\hat{\alpha}_{2}\nu f^{\eta,\hat{\alpha}_{2}}$ as $\phi_{1}$ and setting $f_{in}=\phi_{2}=0$, by Hölder’s inequality, we have
$\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.16)
$\leq C_{\eta}\kappa\left(\hat{\alpha}_{1}\lVert f^{\eta,\hat{\alpha}_{1}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\hat{\alpha}_{2}\lVert f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}\right)$
[…]
This implies that $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$ exists in $L^{2}_{0,\gamma/2}$ for any $0<\eta<1$. Moreover, applying the same procedure as in 3, we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (3.17)
[…]
where $\beta$ will be determined later. Furthermore, we have
$v\cdot\nabla_{x}(f^{\eta})^{2}=-2\frac{1}{\kappa}L(f^{\eta})f^{\eta}+2\phi_{1}f^{\eta}.$ (3.18)
Taking absolute values and integrating over $\Omega\times\mathbb{R}^{3}$ in (3.18), and using (3.17), we deduce that
[…]
Using Lemma 3.1 , for any $\gamma^{\epsilon}$ in (3.1) away from $\partial\Omega_{0}$, we obtain
[…]
By the definition of $P(f)$, we can write $P(f)(x,v)=g(x)m^{1/2}(v)$ for some $g(x)$. When $\epsilon$ is sufficiently small,
$\lVert P(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ $=\int_{\partial\Omega_{2}}|g(x)|^{2}\int_{n(x)\cdot v\geq\varepsilon,\,|v|\leq\tfrac{1}{\varepsilon}}m(v)\,|v\cdot n(x)|\,dv\,dx$
$\geq\int_{\partial\Omega_{2}}|g(x)|^{2}\,dx\times\tfrac{1}{2}\int_{v\cdot n(x)>0}m(v)\,|v\cdot n(x)|\,dv$
$=\tfrac{1}{2}\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (3.21)
Here, we have used the fact that
$\int_{|n(x)\cdot v|\leq\varepsilon}m(v)\,|n(x)\cdot v|\,dv$ $\leq\int_{-\varepsilon^{\prime}}^{\varepsilon}e^{-v_{\parallel}^{2}}|v_{\parallel}|\,dv_{\parallel}\int_{\mathbb{R}^{2}}e^{-|v_{\perp}|^{2}/2}\,dv_{\perp}\;\leq C\varepsilon,$
$\int_{|v|\geq 1/\varepsilon}m(v)\,|n(x)\cdot v|\,dv$ $\leq C\varepsilon.$ (3.22)
We observe that
$\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.23)
$f^{\eta}$ $=P(f^{\eta})+(I-P)(f^{\eta})$ (3.24)
which yields
$\lVert P(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\leq 2\lVert f^{\eta}\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+2\lVert(I-P)(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (3.25)
Combining this estimate with 3.17, 3.20, 3 and 3.25, we conclude that
[…]
Taking $\beta$ sufficiently small and combining 3.17 with 3.26, we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+$ $\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.27)
[…]
Now, similar to the process used to prove that $f^{\hat{\alpha},\eta}$ is a Cauchy sequence for any fixed $\eta$, we observe that
$\left\{\begin{aligned} v\cdot\nabla_{x}\Big(f^{\eta_{1}}-f^{\eta_{2}}\Big)+\frac{1}{\kappa}L\Big(f^{\eta_{1}}-f^{\eta_{2}}\Big)&=0,&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f^{\eta_{1}}-f^{\eta_{2}}&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f^{\eta_{1}}-f^{\eta_{2}}&=\eta_{1}P(f^{\eta_{1}}-f^{\eta_{2}})+\Big(\eta_{1}-\eta_{2}\Big)P(f^{\eta_{2}}),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$.
for any $0<\eta_{1},\eta_{2}<1$. By regarding $\big(\eta_{1}-\eta_{2}\big)P(f^{\eta_{2}})$ as $\phi_{2}$, we obtain
$\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta_{1}}-f^{\eta_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.29)
$\leq C\max\{\kappa,\kappa^{2}\}\Big(\eta_{1}-\eta_{2}\Big)^{2}\lVert P(f^{\eta_{2}})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
[…]
This means that $\{f^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{2}_{0,\gamma/2}$. Therefore $f\coloneqq\lim_{\eta\to 1}f^{\eta}$ is the solution of (3.3). ∎
Chapter 4. $L^{\infty}$ estimate
In this chapter, we denote $w(v)=e^{\alpha|v|^{2}}$. Let $h(x,v)=w(v)f(x,v)$. Then, $h$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}h+\frac{1}{\kappa}L_{w}(h)&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h&=P_{w}(h)+w\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (4.1)
where $P_{w}(\cdot)\coloneq wP(w^{-1}\cdot)$ and $L_{w}(\cdot)\coloneq wL(w^{-1}\cdot)$. Furthermore, we denote
$L_{w}(f)=\nu(v)f-K_{\alpha}(f)$
by defining $K_{\alpha}(\cdot)\coloneq wK(w^{-1}\cdot)$. Note that $K_{\alpha}$ and $|K_{\alpha}|(f)\coloneq w\int_{\mathbb{R}^{3}}|k(v,u)|w^{-1}f(x,u)du$ have the same properties, where $k(v,u)$ denotes the kernel of $K$. In this chapter, we prove that problem (4.1) admits a unique solution in $L^{\infty}$ when $\Omega$ is unsealed with respect to $\partial\Omega_{1}$, and therefore Theorem 1.1 in section 4.2. First, we consider the approximating problem with the parameters $\hat{a}$ and $\eta$, and we define a recursive sequence $h^{l}$ with $h^{1}=0$.
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{l+1}+v\cdot\nabla_{x}h^{l+1}+\frac{1}{\kappa}L_{w}(h^{l+1})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{l+1}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{l+1}&=\eta P_{w}(h^{l})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (4.2)
Second, we establish the following proposition to relate the $L^{\infty}$ norm of $h^{l}$ to the weighted $L^{2}$ norm of $h^{l}/w$. This relation allows us to apply the results from previous chapter to construct a solution to (4.2) in $L^{\infty}$ and then remove the parameters $\hat{a}$ and $\eta$. This approach is known as the $L^{2}-L^{\infty}$ estimate, which is an useful tool ( [13], [14]) for constructing solutions of Boltzmann equation. Finally we note that proposition 4.1 remains valid when $\hat{a}=0$ and $\eta=1$. The statement of the proposition is
Proposition 4.1.
Let $\Omega$ be a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary, and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Suppose that $h^{l+1}$ is the solution of (4.2) and (1.6) holds. Then, there exists $n$ depending only on the constant of unsealed set $\zeta$ such that
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.3)
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{n}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}w^{-1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
To prove this proposition we need the following estimate for $h^{l}$.
$|h^{l+1}|(x,v)$ $\leq\frac{1}{\kappa}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1}|)+S^{\kappa}_{\Omega}(|w\phi_{1}|)+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{1}}|wf_{in}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}|w\phi_{2}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}P_{w}(|h^{l}|).$ (4.4)
After iterating for n times, we obtain
$|h^{l+1}|(x,v)$ $\leq\sum_{k=1}^{4}A_{k}(x,v)+\sum_{j=1}^{n-1}\hat{P}^{j}(A^{j}_{1})(x,v)+\sum_{k=2}^{4}\sum_{j=1}^{n-1}\hat{P}^{j}(A_{k})(x,v)+\hat{P}^{n}(|h^{l+1-n}|)(x,v).$ (4.5)
Here the definitions of $A_{k}^{j}$ and $\hat{P}^{j}$ are given as follows.
  1. 1.
    $A^{j}_{1}(x,v)=\frac{1}{\kappa}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j}|),\quad A_{1}(x,v)=A^{0}_{1}$.
  2. 2.
    $A_{2}(x,v)=S^{\kappa}_{\Omega}(|w\phi_{1}|)$.
  3. 3.
    $A_{3}(x,v)=J^{\kappa}_{1}|wf_{in}|(x,v)=\textbf{1}_{\partial\Omega^{-}_{1}}(q(x,v))e^{-\frac{1}{\kappa}\nu(v)\tau_{-}(x,v)}|w(v)f_{in}|(q(x,v),v)$.
  4. 4.
    $A_{4}(x,v)=J^{\kappa}_{2}|w\phi_{2}|(x,v)=\textbf{1}_{\partial\Omega^{-}_{2}}(q(x,v))e^{-\frac{1}{\kappa}\nu(v)\tau_{-}(x,v)}|w(v)\phi_{2}|(q(x,v),v)$.
  5. 5.
    $\hat{P}(h)(x,v)=J^{\kappa}_{2}wP(w^{-1}h)=J^{\kappa}_{2}\left(W\displaystyle\int hw^{-1}m^{-1/2}d\sigma_{0,1}\right)(x,v)$.
  6. 6.
    $\hat{P}^{j}(h)(x,v)=J^{\kappa}_{2}\left(W\displaystyle\int\cdots\int hw^{-1}m^{-1/2}d\sigma_{j-1,j}d\sigma^{J}_{j-2,j-1}\cdots d\sigma^{J}_{0,1}\right)(x,v)\quad\forall j\geq 2$.
The definitions of the notations appearing above are as follows
  1. 1.
    $W\coloneqq w(v)m^{1/2}(v)$, $x_{0}=q(x,v)$, $x_{j}\coloneqq q(x_{j-1},v_{j})$ for all $j\in\mathbb{N}$.
  2. 2.
    $J^{\kappa}_{i}\coloneqq{1}_{\partial\Omega_{i}}J^{\kappa}$ for $i=1,2$.
  3. 3.
    $\left(\displaystyle\int hd\sigma_{j-1,j}\right)(x_{j-1},v_{j-1})\coloneqq\displaystyle\int_{v_{j}\cdot n(x_{j-1})>0}h(x_{j-1},v_{j})m(v_{j})v_{j}\cdot n(x_{j-1})dv_{j}\quad\forall j\geq 1$.
  4. 4.
    $\left(\displaystyle\int hd\sigma^{1}_{j-1,j}\right)(x_{j-1},v_{j-1})\coloneqq\displaystyle\int_{v_{j}\cdot n(x_{j-1})>0}h(x_{j},v_{j})\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{j})m(v_{j})v_{j}\cdot n(x_{j-1})dv_{j}\,\forall j\geq 1$.
  5. 5.
    $\left(\displaystyle\int hd\sigma^{J}_{j-1,j}\right)(x_{j-1},v_{j-1})\coloneqq\displaystyle\int_{v_{j}\cdot n(x_{j-1})>0}J^{\kappa}_{2}h(x_{j-1},v_{j})m(v_{j})v_{j}\cdot n(x_{j-1})dv_{j}\quad\forall j\geq 1$.
4.1. The existence of solution to (4.1)
We now introduce the definition of an unsealed set:
Definition 4.1.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{2}$ boundary, and let $\partial\Omega_{1}$ be an open subset of $\partial\Omega$. We say that $\Omega$ is unsealed with respect to $\partial\Omega_{1}$ if there exist $0<\zeta<1$ and $m\in\mathbb{N}$ such that
$\displaystyle\int\cdots\int\textbf{1}d\sigma^{1}_{m-1,m}\cdots d\sigma^{1}_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1},$ (4.6)
In particular, $\Omega$ is said to be completely unsealed with respect to $\partial\Omega_{1}$ when $m=1$.
Remark 4.1.
The $m$ can be replaced by $m_{T}$ in Definition 4.1 . Additional information about unsealed set is given in Appendix 6.
Then, we prove proposition 4.1 and state our main goal of this section. The statement is as follows:
Theorem 4.1.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary and suppose $\Omega$ is unsealed with respect to $\partial\Omega_{1}$. Assume that $3<\gamma\leq 1$, $0\leq a<\frac{1}{4}$ and (1.6) holds. Then, for any $\phi_{1}\in L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3}),f_{in}\in L^{\infty}_{\alpha,0}(\partial\Omega_{1}),$ and $\phi_{2}\in L^{\infty}_{\alpha,0}(\partial\Omega_{2})$, there exists a unique solution $h\in L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$ to (4.1). Furthermore. the solution $h$ satisfies following estimate
$\lVert h\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\lVert h\rVert_{L^{\infty}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big).$ (4.7)
For clarity, we present the proof for the case of $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. The proof for unsealed case is similar. In order to achieve our goal, we need the following lemmas to establish proposition 4.1.
Lemma 4.1.
Let $3<\gamma\leq 1$ and assume that (1.6) holds. Then
$|h^{l+1}|(x,v)$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.8)
$+\frac{\zeta^{n-1}}{1-\zeta}\max_{0\leq j\leq n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\left(\frac{1}{\kappa}\right)^{2}\sum_{j=0}^{n-1}\sum_{i=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|)(x,v)$
for any $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$.
Proof.
For $A_{3}$ and $A_{4}$, because $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$, we have
$\begin{cases}\lVert A_{3}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert A_{4}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert w(v)\phi_{2}\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.\end{cases}$ (4.9)
Also, similar to (4.9), we obtain
$\begin{cases}\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{3})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{4})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}.\end{cases}$ (4.10)
Using the Proposition 2.2, we obtain
$\lVert A_{2}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\kappa\lVert\langle v\rangle^{-1}w(v)\phi_{1}\rVert_{L^{\infty}(\partial\Omega^{-}_{1})}$ (4.11)
and
$\left\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{2})\right\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\frac{\kappa}{1-\zeta}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}.$ (4.12)
Applying that $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$, we again yield
$\lVert\hat{P}^{n}(h^{l+1-n})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\zeta^{n-1}\lVert h^{l+1-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}.$ (4.13)
Combine the above result, we conclude that
$|h^{l+1}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.14)
$\quad+\frac{1}{\kappa}\sum_{j=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j}|)+\zeta^{n-1}\lVert h^{l+1-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}.$
For each 0<j<n10<j<n-1, replace $l+1$ to l+1jl+1-j, we have
$|h^{l+1-j}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.15)
$\quad+\frac{1}{\kappa}\sum_{i=0}^{n-1}\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|)+\zeta^{n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}.$
Substitute $h^{l+1-j}$ by 4.15, we obtain
$|h^{l+1}|$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.16)
$\quad+\frac{1}{\kappa}\sum_{j=0}^{n-1}\zeta^{j}\left\lVert\displaystyle S^{\kappa}_{\Omega}|K_{\alpha}|\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big)\right\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+\frac{1}{\kappa}\sum_{j=0}^{n-1}\zeta^{j}\left\lVert\displaystyle S^{\kappa}_{\Omega}|K_{\alpha}|\Big(\kappa\lVert\phi_{1}\rVert_{L_{\alpha,-1}^{\infty}(\Omega\times\mathbb{R}^{3})}+\zeta^{n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\Big)\right\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+\left(\frac{1}{\kappa}\right)^{2}\sum_{j=0}^{n-1}\sum_{i=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|)$
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+\frac{\zeta^{n-1}}{1-\zeta}\max_{0\leq j\leq n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\left(\frac{1}{\kappa}\right)^{2}\sum_{j=0}^{n-1}\sum_{i=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|).$ (4.17)
For convenience, we define
$A_{ij}(f)\coloneq\left(\frac{1}{\kappa}\right)^{2}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(f)$ (4.18)
and denote $\mathcal{D}=\text{diam}(\Omega)$. Now, we prove the following lemma
Lemma 4.2.
Suppose $\Omega$ is a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Given $\epsilon>0$ and $i,j\in\mathbb{N}$. Suppose $3<\gamma\leq 1$ and (1.6) holds. Then
$\lVert A_{ij}(f)\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\leq C\zeta^{j+i-2}\Big(\epsilon\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.19)
Proof.
Without loss of generality, we assume $f\geq 0$. For $i=j=0$, we use properties of $|k_{\alpha}|$ like 2.3, proposition 2.6 , and proposition 2.5. By definition,
$A_{00}$ $=\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{\mathbb{R}^{3}}\int_{0}^{\tau_{-}(x-tv,u)}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (4.20)
Fixed some $\epsilon_{1},\epsilon_{2}>0$ and $N,M\in\mathbb{N}$, we define
$A=\{|v-u|>\epsilon_{2},|u|<N\}\text{ and }B=\{|u-w|>\epsilon_{1}\}.$ (4.21)
Moreover, we assume that $f\geq 0$. Then, we have
$A_{00}$ (4.22)
$\leq C\epsilon^{3-|\gamma|}_{1}\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{\mathbb{R}^{3}}\int_{0}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{\{|v-u|>\epsilon_{2}\}}\int_{0}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{A}\int_{0}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{A}\int_{1/N}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (4.23)
Using Hölder inequality and noticing that $k(v,u),k(u,w)$ are bounded on $A,B$ respectively. Let y=xtvsuy=x-tv-su, we obtain
$A_{00}$ $\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{1+M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$+C\textbf{1}_{|v|<M}\left(\frac{1}{\kappa}\right)^{2}\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{A}\int_{1/N}^{\infty}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+C\left(\frac{1}{\kappa}\right)^{2}\textbf{1}_{|v|<M}\Big(\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{\mathbb{R}^{3}}\int_{1/N}^{\infty}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v)t}\frac{1}{s^{3}}\nu(w)f^{2}(y,w)w^{-2}(w)dwdsdydt\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\left(\frac{1}{\kappa}\right)\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (4.24)
Hence, if $\epsilon_{1},\epsilon_{2}$ are sufficiently small and $N,M$ are chosen large enough, the proof for $A_{00}$ is complete. For $i$ or $j\neq 0$, since $wm^{1/2}\leq C$ and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$, it suffice to estimate $S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(f)$. Furthermore, we can observe that the character of $|K_{\alpha}|$ can be replace by $\hat{P}^{i}$. Therefore, the estimate $A_{ij}$ is similar to $A_{00}$ and it is sufficient to estimate $\frac{1}{\kappa}\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(f)$. For instance, let $i=1$ and denote $q=q(x,v)$,
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (4.25)
$=\frac{1}{\kappa}\displaystyle\int W(v)J^{-}_{2}(q(x,v),v)S^{\kappa}_{\Omega}|K_{\alpha}|(f)(q(x,v),v_{1})w^{-1}(v_{1})d\sigma_{0,1}$
$\leq\frac{1}{\kappa}\displaystyle\int_{v_{1}\cdot n(q)>0}\int_{0}^{\tau_{-}(q,v_{1})}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k_{\alpha}|(v_{1},u)f(q-tv_{1},u)w^{-1}(v_{1})m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}$
$=\frac{1}{\kappa}\displaystyle\int_{v_{1}\cdot n(q)>0}\int_{0}^{\tau_{-}(q,v_{1})}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}.$
Similar to $A_{00}$, fix some $\epsilon_{1}>0$ and $N\in\mathbb{N}$, we define
$A=\{|v_{1}|<N\},\quad B=\{|v_{1}-u|>\epsilon_{1}\}.$ (4.26)
Then,
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (4.27)
$\leq C\epsilon^{3-|\gamma|}_{1}\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+\frac{1}{\kappa}\displaystyle\int_{v_{1}\cdot n(q)>0}\int_{0}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+\frac{1}{\kappa}\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{0}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}.$ (4.28)
Using Hölder inequality and notice that $k(v_{1},u)$ is bounded on $B$. Let $y=q(x,v)-tv_{1}$. We obtain
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.29)
$\quad+\frac{1}{\kappa}\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{\frac{\mathcal{D}}{N}}^{\mathcal{D}/|v_{1}|}\int_{B}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+C\frac{1}{\kappa}\Big(\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{\frac{\mathcal{D}}{N}}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}|k|(v_{1},u)f^{2}(q-tv_{1},u)w^{-2}(u)dudtdv_{1}\Big)^{1/2}$ (4.30)
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+C\frac{1}{\kappa}\Big(\displaystyle\int_{\Omega}\int_{\frac{\mathcal{D}}{N}}^{\infty}\int_{\mathbb{R}^{3}}\frac{1}{t^{3}}\nu(u)f^{2}(y,u)w^{-2}(u)dudtdy\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\left(\frac{1}{\kappa}\right)\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (4.31)
Hence, when $\epsilon_{1}$ is sufficiently small and $N$ is sufficiently large, the proof for $A_{ij}$ is complete. ∎
Proof.
The proof of property 4.1
We use Lemma 4.1 and Lemma 4.2 to obtain
$|h^{l+1}|(x,v)$ (4.32)
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+\frac{\zeta^{n-1}}{1-\zeta}\max_{0\leq j\leq n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$
$\quad+\frac{C}{\zeta^{2}(1-\zeta)^{2}}\Big(\epsilon\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
Taking $n$ is enough large and $\epsilon$ is sufficiently small, we have
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.33)
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
The proof is complete. ∎
Proof.
The proof of Theorem 4.1
Now, we construct the solution of (4.1) by starting from (4.2). It is clear that Property 4.1 still holds when $\hat{\alpha}=0$ or $\eta=1$. We rewrite 4.33 as
$\lVert h^{i+2(n-1)+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.34)
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.35)
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\left(\min\{1,\frac{1}{\kappa}\}\right)\max_{0\leq m\leq 2(n-1)}\lVert h^{i+1+m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})},$
for any $i=0,1,\cdots$. By Lemma 4.3 with $a_{i}=\lVert h^{i}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ and k=2(n1)k=2(n-1) and notice that
$\lVert h^{i}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ $=\lVert f^{i}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (4.36)
$\leq C_{\eta}\max\{\sqrt{\kappa},\kappa\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
for any $i\in\mathbb{N}$. Then
$\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{i}{2n-1}]}\max_{0\leq m\leq 4(n-1)}\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.37)
$\quad+C_{n,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+C_{\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
The statement of Lemma 4.3 is
Lemma 4.3 ([13](lemma 6.1)).
Let $\{a_{i}\}_{i=0}\in\mathbb{R}_{+}$. For any fixed $k\in\mathbb{N}$, we denote $A^{k}_{i}=\max\{a_{i},a_{i+1},\cdots,a_{i+k}\}$.
  1. 1.
    Assume $D\geq 0$. If $a_{i+1+k}\leq\frac{1}{8}A_{i}^{k}+D$ for $i=0,1,\cdots,$ then it holds that
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (4.38)
  2. 2.
    Let $0\leq\eta<1$ with $\eta^{k+1}\geq\frac{1}{4}$. If $a_{i+1+k}\leq\frac{1}{8}A_{i}^{k}+C_{k}\eta^{i+k+1}$ for $i=0,1,\cdots$, then it holds that
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+2C_{k}\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (4.39)
The proof of Lemma 4.3 is omitted as it appears in the appendix of [13]. Now, we should control $\max_{0\leq m\leq 4(n-1)}\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$. We can use same techniques in the proof of Property 4.1 and induction to show that
$\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{n,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.40)
$\quad+C_{n,\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
for $0\leq m\leq 4(n-1)$. Here, $C_{n,\eta},C_{n,\zeta}$ are increasing with respect to $n$. In particular, combining 4.37 and 4.40, we have
$\lVert h^{l}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{l}{2n-1}]}C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.41)
$\qquad+C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big),$ (4.42)
since $L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})\subset L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$.
Now, we show that $\{h^{l}\}$ is a Cauchy sequence in $L^{\infty}(\Omega\times\mathbb{R}^{3})$. Consider $2n\ll m_{1},m_{2}$, we notice that $h^{m_{1}+1}-h^{m_{2}+1}$ satisfies
$\left\{\begin{aligned} &\hat{\alpha}\nu(v)\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+v\cdot\nabla_{x}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+\frac{1}{\kappa}L_{w}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=\eta P_{w}\left(h^{m_{1}}-h^{m_{2}}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
By the same procedure as in Property 4.1, we obtain
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.43)
$\qquad\qquad+C_{n}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
Note that $\{f_{m}\}$ converges in $L^{2}_{0,\gamma/2}$ (see the proof of theorem 3.1). Given $\epsilon>0$, by definition, there exists $M$ such that for any $m_{1},m_{2}>M$, we have
$C_{n}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}<\frac{\epsilon}{2(2n+7)/7}.$ (4.44)
we then obtain
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\epsilon.$ (4.45)
By (4.41), we find that
$\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{f_{in},\phi_{1},\phi_{2}}C_{n,\eta,\zeta}\left(\frac{1}{8}\right)^{[\frac{\min\{m_{1},m_{2}\}}{2n-1}]}+\epsilon/2.$ (4.46)
Hence, if $m_{1},m_{2}$ satisfy
$C_{f_{in},\phi_{1},\phi_{2}}C_{n,\eta,\zeta}\left(\frac{1}{8}\right)^{[\frac{\min\{m_{1},m_{2}\}}{2n-1}]}<\epsilon/2,$
then it follows that $\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}<\epsilon$. From the discussion above, $\{h^{l}\}$ is a Cauchy sequence. Therefore, the limit $\lim_{l\to\infty}h^{l}=h^{\hat{\alpha},\eta}$ exists for any $\alpha,\eta$. Moreover, $h^{\hat{\alpha},\eta}$ satisfies
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{\hat{\alpha},\eta}+v\cdot\nabla_{x}h^{\hat{\alpha},\eta}+\frac{1}{\kappa}L_{w}(h^{\hat{\alpha},\eta})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{\hat{\alpha},\eta}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{\hat{\alpha},\eta}&=\eta P_{w}(h^{\hat{\alpha},\eta})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Moreover, we observe that the argument used in Property 4.1 also applies to $h^{\hat{\alpha},\eta}$ for any fixed $0<\eta<1$. Then we obtain
$\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.47)
$\qquad+C_{n}\left(\frac{1}{\kappa}\right)\lVert h^{\hat{\alpha},\eta}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$=C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\qquad+C_{n,\eta}\left(\frac{1}{\kappa}\right)\lVert f^{\hat{\alpha},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
Let $0<\hat{\alpha}_{1},\hat{\alpha}_{2}<1$. We recall that
$\left\{\begin{aligned} v\cdot\nabla_{x}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)+\frac{1}{\kappa}L_{w}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=-\hat{\alpha}_{1}h^{\hat{\alpha}_{1},\eta}+\hat{\alpha}_{2}h^{\hat{\alpha}_{2},\eta},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ \left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ \left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=\eta P_{w}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Regard $\hat{\alpha}_{1}h^{\hat{\alpha}_{1},\eta}+\hat{\alpha}_{2}h^{\hat{\alpha}_{2},\eta}$ as $\phi_{1}$. By repeating the same process as above, we obtain
$\lVert h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.48)
$\leq C_{\zeta}\kappa\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert\langle v\rangle^{-1}h^{\hat{\alpha}_{1},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\langle v\rangle^{-1}h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{n,\eta}\min\{1,\frac{1}{\kappa}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\kappa\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{n,\eta}\min\{1,\frac{1}{\kappa}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
This implies that $\{h^{\hat{\alpha},\eta}\}$ is a Cauchy sequence in $L^{\infty}(\Omega\times\mathbb{R}^{3})$. After we denote $h^{\eta}=\displaystyle\lim_{\alpha\to 0}h^{\hat{\alpha},\eta}$, note that $h^{\eta_{1}}-h^{\eta_{2}}$ satisfies
$\left\{\begin{aligned} &v\cdot\nabla_{x}\left(h^{\eta_{1}}-h^{\eta_{2}}\right)+\frac{1}{\kappa}L_{w}\left(h^{\eta_{1}}-h^{\eta_{2}}\right)=0,&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ &\left(h^{\eta_{1}}-h^{\eta_{2}}\right)=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ &\left(h^{\eta_{1}}-h^{\eta_{2}}\right)=\eta_{1}P_{w}\left(h^{\eta_{1}}-h^{\eta_{2}}\right)+\left(\eta_{1}-\eta_{2}\right)P_{w}\left(h^{\eta_{2}}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
We can regard $\left(\eta_{1}-\eta_{2}\right)P_{w}\left(h^{\eta_{2}}\right)$ as $\phi_{2}$ and apply the same argument again. We obtain
$\lVert h^{\eta_{1}}-h^{\eta_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert P_{w}(h^{\eta_{2}})\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}+C_{n}\left(\frac{1}{\kappa}\right)\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (4.49)
$\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert h^{\eta_{2}}\rVert_{L^{\infty}(\partial\Omega^{+})}+C_{n}\left(\frac{1}{\kappa}\right)\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\qquad+C_{n}\min\{1,\frac{1}{\kappa}\}\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
The last inequality follows from a similar argument showing that $\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}}$ is bounded with respect to $\hat{\alpha},\eta$. Thus $\{h^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{\infty}$ and $h\coloneq\displaystyle\lim_{\eta\to 1}h^{\eta}$ is a mild solution of (4.1). The proof of the estimate (4.7) is similar to the proof of (4.47). ∎
Remark 4.2.
the proof for the unseal set is similar to the case of the complete unsealed set. We just take $\frac{n}{m}]$ sufficiently large in the proof of Proposition 4.1 to make $\zeta^{[\frac{n}{m}]}$ is small enough.
4.2. the proof of Theorem 1.1
Now, we conclude this chapter by proving Theorem 1.1.
Proof.
the proof of theorem 1.1 We consider the following iteration with $f_{1}=0$.
$\left\{\begin{aligned} v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})=&\frac{1}{\kappa}\Gamma(f_{k},f_{k}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f_{k+1}=&f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}=&P(f_{k+1})+m^{-1}(m_{T(x)}-m)P(f_{k})\\ &\quad+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (4.50)
By 4.1, we obtain
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&\lVert m^{-1}(m_{T(x)}-m)P(f_{k})\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\frac{1}{\kappa}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$
$\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&C\lVert m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\lVert f_{k}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\frac{1}{\kappa}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$. (4.51)
Let $\mathcal{A}=\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}$ and $\mathcal{B}=\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}$. It is well-known that
$\lVert\Gamma(f,f)\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\leq C\lVert f\rVert^{2}_{L^{\infty}_{a,0}(\Omega\times\mathbb{R}^{3})}.$ (4.52)
and if $0<\alpha<1/4$ and $T(x)<2$, then
$\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\leq C_{\alpha}\lVert T(x)-1\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.$ (4.53)
We show that there exists $\delta$ such that, if $\mathcal{A}+\mathcal{B}\leq\delta$, then
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}(\mathcal{A}+\mathcal{B})$ (4.54)
for all $k\in\mathbb{N}$. The constant $C_{\zeta}$ is defined in Theorem 4.1.
For $k=1$,
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C_{\zeta}(\mathcal{A}+\mathcal{B})$
by (4.2). Suppose the statement holds for $k=n$ . For $k=n+1$,
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (4.55)
$\leq C_{\zeta}(\mathcal{A}+\mathcal{B})+C_{\zeta}C\mathcal{B}(\mathcal{A}+\mathcal{B})+4CC^{2}_{\zeta}\frac{1}{\kappa}(\mathcal{A}+\mathcal{B})^{2}$ (4.56)
$\leq\Big(C_{\zeta}+C_{\zeta}C\mathcal{B}+4CC^{2}_{\zeta}\max\{1,\frac{1}{\kappa}\}(\mathcal{A}+\mathcal{B})\Big)(\mathcal{A}+\mathcal{B}).$
Choose $\delta<\min\{\frac{1}{4CC_{\zeta}},\frac{\min\{1,\kappa\}}{32CC^{2}_{\zeta}}\}$. Then, the proof is complete. Now, consider $f_{k+1}-f_{k}$ for $k\in\mathbb{N}$. Following a similar procedure, we obtain
[…]
This implies that $\{f_{k}\}$ is a Cauchy sequence in $L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$. Hence, we obtain the solution by taking the limit $f=\displaystyle\lim_{k\to\infty}f_{k}$. Uniqueness follows by same argument. Further the upper bound of $f$, by (4.55), we have
$\lVert f\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C(\mathcal{A}+\mathcal{B})\leq C\delta.$ (4.57)
Chapter 5. Application to Thermal transpiration and thermomolecular pressure difference
5.1. the model of the transpiration : the proof of Theorem 1.2 and Theorem 1.3
In this chapter, we prove the Theorem 1.2. Without loss of generality, we assume that $a=0$ and $b=L$. To simplify the calculation and to obtain a more explicit estimate, we prove Theorem 1.2 when $\Omega$ satisfies
$\Omega\cap\{0\leq x_{3}\leq L\}$ $=\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}|\,x_{1}^{2}+x_{2}^{2}\leq 1,0\leq x_{3}\leq L\}.$
The boundary condition is imposed as follows.
$\begin{aligned} F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{v^{2}}{2}},&&\,\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}\cap\{x_{3}<0\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{v^{2}}{2T_{2}}},&&\,\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}\cap\{x_{3}>L\},\\ F(x,v)&=m_{T(x)}\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\,\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}$ (5.1)
Without loss of generality, we assume that $T_{2}>1$. We rewrite
$F(x,v)=m_{T_{q}}(x,v)+f(x,v)=\frac{1}{(2\pi)^{3/2}T^{2}(q(x,v))}e^{-\frac{|v|^{2}}{2T(q(x,v))}}+f(x,v).$ (5.2)
Recall that $m_{T_{q}}(x,v)$ is a solution of
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=0,&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (5.3)
The proof appears in the [36]. Now, the equation (1.5) becomes
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=\frac{1}{\sqrt{2\pi}}\left(\frac{1}{\sqrt{T_{2}}}-1\right)\frac{1}{2\pi T_{2}^{2}}e^{-\frac{|v|^{2}}{2T_{2}}}\coloneq-C_{T_{2}}m_{T_{2}},&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{(2\pi)T^{2}(x)}e^{-\frac{|v|^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (5.4)
Note that $C_{T_{2}}>0$. Recall that the local flux $u(x)$ and the total flux $U(z)$ in the $x_{3}$-direction are defined as
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=z\}}u(x)dS.$ (5.5)
Now, we establish a lemma concerning $\Omega$ and $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$.
Lemma 5.1.
$\Omega$ is completely unsealed with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$.
Proof.
For any fixed $x\in\overline{\Omega\cap\{0\leq x_{3}\leq L\}}$, we classify $v\in\mathbb{R}^{3}$ as follows.
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\partial\Omega_{1,a}^{-}\right\},$ (5.6)
$A_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\partial\Omega_{1,b}^{-}\right\},$
$A_{3}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\partial\Omega_{2}^{-}\right\},$
Let $v=(v_{1},v_{2},v_{3})=(\rho\cos\phi\sin\theta,\rho\sin\phi\sin\theta,\rho\cos\theta)$ and $\hat{v}=(v_{1},v_{2}),\,\hat{x}=(x_{1},x_{2})$. Then the sets $A_{i}$ can be written more explicitly as follow.
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}<\cos\theta\leq 1\right\},$ (5.7)
$A_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,-1\leq\cos\theta<\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\right\},$
$A_{3}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\leq\cos\theta\leq\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}\right\},$
where
$r\coloneq h+\sqrt{h^{2}+k^{2}},\quad h\coloneq\hat{x}\cdot\frac{\hat{v}}{|\hat{v}|},\quad k\coloneq\sqrt{1-|\hat{x}|^{2}}.$
Recalling Remark 4.1, we now consider the case $x\in\partial\Omega_{2}^{-}$. Without loss of generality, we assume that $(x_{1},x_{2})=(1,0)$. Then,
$\int_{v\cdot n(x)>0}\textbf{1}_{\partial\Omega_{2}}\circ q(x,v)m_{T}(x,v)v\cdot n(x)dv$ $=2\int_{-\pi/2}^{\pi/2}\int_{\cos^{-1}\left(\frac{x_{3}}{\sqrt{x_{3}^{2}+4\sin^{2}\phi}}\right)}^{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+4\sin^{2}\phi}}\right)}\cos\phi\sin^{2}\theta d\theta d\phi$
$\coloneq\Theta(x_{3}).$
Observe that $0\leq\Theta(x_{3})\leq 1$ and $\Theta(x_{3})$ is a continuous on $[0,L]$. Therefore, there exists $z_{0}\in[0,L]$ such that $\max_{[0,L]}\Theta(x_{3})=\Theta(z_{0})$. However, $\Theta(z_{0})$ is equal to $1$ if and only if
$\cos^{-1}\left(\frac{z_{0}-L}{\sqrt{(z_{0}-L)^{2}+4\sin^{2}\phi}}\right)=\pi\quad\text{and}\quad\cos^{-1}\left(\frac{z_{0}}{\sqrt{z_{0}^{2}+4\sin^{2}\phi}}\right)=0$ (5.8)
for all $\phi$. This is impossible. Hence,
$\int_{v\cdot n(x)>0}\textbf{1}_{\partial\Omega_{2}}\circ q(x,v)m_{T}(x,v)v\cdot n(x)dv$ $\leq\Theta(z_{0})<1\quad\text{for all}\,x\in\partial\Omega_{2}.$
Note that if $T_{2}$ is sufficiently close to $1$ and $\Omega$ has $C^{2}$ boundary, then Theorem 1.1 guarantees the existence of $F$ in $L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})$.
Proof.
The proof of theorem 1.2 for the pipe case.
Note that the $f$ in 5.4 has the form
$f(x,v)$ $=-\textbf{1}_{\partial\Omega^{-}_{1,b}}(q(x,v))C_{T_{2}}m_{T_{2}}$ (5.9)
$\quad+\textbf{1}_{\partial\Omega^{-}_{2}}(q(x,v))m_{T}(q(x,v),v)\int_{v\cdot n(x)>0}f(q(x,v),v)v\cdot n(x)dv$
$\quad+\frac{1}{\kappa}\left(\int_{0}^{\tau_{-}(x,v)}\left[2Q(m_{T_{q}},f)+Q(f,f)+Q(m_{T_{q}},m_{T_{q}})\right](x-sv,v)ds\right)$
$\coloneq\textbf{1}_{\partial\Omega^{-}_{1,b}}g(x,v)+J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(f)(x,v)+\frac{1}{\kappa}h[f](x,v).$
Observe that
$\int_{\mathbb{R}^{3}}v_{3}m_{T_{q}}dv=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{e^{-\frac{\rho^{2}}{2T(\theta,\phi)}}}{2\pi T^{2}(\theta,\phi)}\rho^{3}\cos\theta\sin\theta d\theta d\phi d\rho=0.$
We now use a procedure similar to the one in Chapter 4 . When tracing back $n+1$ time for $n\geq 1$, we obtain
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (5.10)
$=\int_{\mathbb{R}^{3}}v_{3}\textbf{1}_{\partial\Omega^{-}_{1,b}}gdv+\frac{1}{\kappa}\int_{\mathbb{R}^{3}}v_{3}h[f]dv+\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(f)dv$
$=\int_{\mathbb{R}^{3}}v_{3}\textbf{1}_{\partial\Omega^{-}_{1,b}}gdv+\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)dv$
$\quad+\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)+\frac{1}{\kappa}\int_{\mathbb{R}^{3}}v_{3}h[f]dv$
$\quad+\frac{1}{\kappa}\int_{\mathbb{R}^{3}}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(h[f])dv+\frac{1}{\kappa}\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(h[f])dv$
$\quad+\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{n+1}\circ\overline{P}(f)dv$
$\coloneq A_{1}+A_{2}+\sum_{j=1}^{n}B_{j}+\frac{1}{\kappa}\Big(C_{1}+C_{2}\Big)+D.$
The definitions of $\overline{P}(f)$, $\hat{P}^{j}(f)$ and $J^{T}_{\partial A}$ are as follows.
$J^{T}_{A}(f)(x,v)$ $\coloneq\textbf{1}_{A}(q(x,v))m_{T}(q(x,v),v)f(q(x,v)),$
$\overline{P}(f)(x,v)$ $\coloneq\Big(\int_{v_{0}\cdot n(x)>0}f(x,v_{0})v_{0}\cdot n(x)dv_{0}\Big)(x,v),$
$\hat{P}^{j}(f)(x,v)$ $\coloneq\Big(\int\cdots\int fd\sigma^{T}_{j-1,j}d\sigma^{T}_{j-2,j-1}\cdots d\sigma^{T}_{0,1}\Big)(x,v),$
$\Big(\displaystyle\int fd\sigma^{T}_{j-1,j}\Big)(x_{j-1},v_{j-1})$ $\coloneqq\displaystyle\int_{v_{j}\cdot n(x_{j-1})>0}J^{T}_{\partial\Omega^{-}_{2}}(f)(x_{j-1},v_{j})v_{j}\cdot n(x_{j-1})dv_{j}.$
Here $x_{0}=q(x,v)$, $v_{0}=v$, and $x_{i}=q(x_{i-1},v_{i})$ for all $i\in\mathbb{N}$.
Now, we estimate each term in (5.10). The definition of $A_{1}$ to $D$ appearing in the last line of (5.10) are:
$A_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}\textbf{1}_{\partial\Omega^{-}_{1,b}}gdv,$ (5.11)
$A_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)dv,$ (5.12)
$B_{j}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g),$ (5.13)
$C_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}h[f]dv,$ (5.14)
$C_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(h[f])dv+\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(h[f])dv,$ (5.15)
$D$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{n+1}\circ\overline{P}(f)dv.$ (5.16)
For $A_{1}$, using (5.7). we obtain
$A_{1}$ $=-\int_{0}^{\infty}\int_{0}^{2\pi}\int_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\right)}^{\pi}C_{T_{2}}\frac{e^{-\frac{\rho^{2}}{2T_{2}}}}{2\pi T_{2}}\rho^{3}\cos\theta\sin\theta d\theta d\phi d\rho$
$=C\left(\int_{0}^{2\pi}\frac{r^{2}}{(x_{3}-L)^{2}+r^{2}}d\phi\right)\left(1-\frac{1}{\sqrt{T_{2}}}\right)$
$>0.$
For $A_{2}$, we observe that
$\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)(x)$ $=\int_{v_{0}\cdot n(x)>0}\textbf{1}_{\partial\Omega^{-}_{1,b}}gv_{0}\cdot n(x)dv_{0},$
$=2C\int_{-\pi/2}^{\pi/2}\int^{\pi}_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+4\sin^{2}\phi}}\right)}\cos\phi\sin^{2}\theta d\theta d\phi$
$\coloneq-\Theta(x_{3})\left(1-\frac{1}{\sqrt{T_{2}}}\right).$
Let $v=(\rho\cos\phi\sin\theta,\rho\sin\phi\sin\theta,\rho\cos\theta)$. Then, we have
$A_{2}$ $=\frac{1}{T_{2}^{2}}\left(1-\frac{1}{\sqrt{T_{2}}}\right)\int_{0}^{\infty}\int_{0}^{2\pi}\int_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\right)}^{\cos^{-1}\left(\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}\right)}m_{T}\Theta(q(x,v))\rho^{3}\cos\theta\sin\theta d\theta d\phi d\rho$
$\coloneq\overline{\Theta}_{1}(x)\left(1-\frac{1}{\sqrt{T_{2}}}\right).$
Similarly, there exists $\overline{\Theta}_{2,n}(x)$ such that
$\sum_{j=1}^{n}B_{j}=\overline{\Theta}_{2,n}(x)\left(1-\frac{1}{\sqrt{T_{2}}}\right).$
The relation between the explicit forms of $\Theta(x)$, $\overline{\Theta}_{1}(x)$ and $\overline{\Theta}_{2}(x)$ is given by
$\Theta(x)$ $=2\int_{-\pi/2}^{\pi/2}\int^{\pi}_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+4\sin^{2}\phi}}\right)}\cos\phi\sin^{2}\theta d\theta d\phi,$ (5.17)
$\overline{\Theta}_{1}(x)$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\Theta dv$ (5.18)
$=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\right)}^{\cos^{-1}\left(\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}\right)}m_{T}\Theta(q(x,v))\rho^{3}\cos\theta\sin\theta d\theta d\phi d\rho,$
$\overline{\Theta}_{2}(x)$ $=\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}(\Theta(x)).$ (5.19)
$=\sum_{j=1}^{n}\int_{0}^{\infty}\int_{0}^{2\pi}\int_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+r^{2}}}\right)}^{\cos^{-1}\left(\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}\right)}m_{T}\hat{P}^{j}(\Theta)(q(x,v))\rho^{3}\cos\theta\sin\theta d\theta d\phi d\rho,$
where $v=(\rho\cos\phi\sin\theta,\rho\sin\phi\sin\theta,\rho\cos\theta)$. Since $\Omega$ is completely unsealed with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$ and applying Proposition 2.9 , we obtain
$\frac{1}{\kappa}\Big(|C_{1}|+|C_{2}|\Big)\leq\frac{C_{\zeta}}{\kappa}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})},\quad|D|\leq C\zeta^{n+1}\lVert f\rVert_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}.$
Observe that $\overline{\Theta}_{1}(x)$ and $\overline{\Theta}_{2,n}(x)$ are positive for any $x\in\{x_{3}=0\}$. After combining all estimates and taking $n$ and $\kappa$ sufficiently large, there exists a constant $C$ such that
$U(0)=\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=0\}}u(x)dS\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (5.20)
Then, we finish the proof by invoking the conservation law of mass. ∎
For the general case, the terms $A_{1},A_{2}$ and $B_{j}$ are very difficult to write down in explicit form, since (5.7) depends on $\partial\Omega\cap\{x_{3}=0\}$ and $\partial\Omega\cap\{x_{3}=L\}$. However, we can still prove that $\Omega$ is a completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$, and the proof of the general case proceed in the same way as before.
Lemma 5.2.
$\Omega$ is completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$
Proof.
Since $\Omega$ is an open set, there exist $\rho_{1},\rho_{2}>0$ and $a_{1},a_{2}\in\mathbb{R}^{3}$ such that
$B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\subset\Omega\cap\{x_{3}=0\}\quad\text{and}\quad B(a_{2},2\rho_{2})\cap\{x_{3}=L\}\subset\Omega\cap\{x_{3}=L\}.$ (5.21)
Without loss of generality, we may assume that $a_{1}=0$ and check the definition 4.1 when $x\in\partial\Omega_{2}$ such that $x=(x_{1},0,x_{3})$ with $x_{1}>\rho_{1}$. We classify $v\in\mathbb{R}^{3}$ as follows
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\left[B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\right]^{-}\right\},$
$A_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\left[B(a_{2},2\rho_{2})\cap\{x_{3}=0\}\right]^{-}\right\},$
and
$\mathcal{A}_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}}{x_{1}+\rho_{1}}<\tan\theta<\frac{x_{3}}{x_{1}-\rho_{1}},\,|\tan\phi|<\frac{h_{1}}{x_{3}},0\leq\phi\leq\pi\right\},$ (5.22)
$\mathcal{A}_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{L-x_{3}}{d+\rho_{2}}<\tan\theta_{1}<\frac{L-x_{3}}{d-\rho_{2}},\,|\tan\phi_{2}|<\frac{h_{2}}{L-x_{3}},\pi\leq\phi_{2}\leq 2\pi\right\},$
$\mathcal{A}_{3}$ $=\mathbb{R}^{3}\setminus\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right),$
where
$d$ $=\text{distance between}\,(x_{1},0,L)\,\text{and}\,a_{2},$
$\hat{r}_{1}$ $\coloneq x_{1}-x_{3}\cot\theta,\quad\hat{r}_{2}\coloneq d-(L-x_{3})\cot\theta,$
$h_{1}$ $\coloneq\sqrt{\rho^{2}_{1}-\hat{r}^{2}},\quad h_{2}\coloneq\sqrt{\rho^{2}_{2}-\hat{r}_{2}^{2}}.$
and $v=(v_{1},v_{2},v_{3})=(\rho\cos\theta,\rho\cos\phi\sin\theta,\rho\sin\phi\sin\theta)$ respect to standard basis for $\mathcal{A}_{1}$ but $v=(v_{1},v_{2},v_{3})=(\rho\cos\theta_{1},\rho\cos\phi_{1}\sin\theta_{1},\rho\sin\phi_{1}\sin\theta_{1})$ respect to the basis as
$e_{1}=\frac{x-a_{2}}{|x-a_{2}|},\quad\{e_{1},e_{2},e_{3}\}\,\text{is an orthogonal basis of }\mathbb{R}^{3}$
for $\mathcal{A}_{2}$. Also,
$\begin{aligned} \mathcal{A}_{1}&=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,0\leq\phi\leq\pi\right\}\quad&&\text{on }x_{3}=0,\\ \mathcal{A}_{2}&=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\pi\leq\phi\leq 2\pi\right\}\quad&&\text{on }x_{3}=L.\end{aligned}$
Note that $\mathcal{A}_{i}\subset A_{i}$ for any $i=1,2$. Define
$\partial\mathcal{D}$ $\coloneq\partial\left(\Omega\cap\{0\leq x_{3}\leq L\}\right)\setminus\Big\{\big(B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\big)\cup\big(B(a_{2},2\rho_{2})\cap\{x_{3}=L\}\big)\Big\},$
$\partial\mathcal{A}_{3}$ $\coloneq\{(x,v)\in\partial\mathcal{D}\times\mathbb{R}^{3}|\,n(x)\cdot v<0\}.$
Notice that $\partial\mathcal{D}$ is not $C^{2}$, so the definition of $n(x)$ is taken from the original charts of $\partial\Omega$.
Now, since $\partial\mathcal{A}_{3}$ is a compact set in $\mathbb{R}^{3}$ and $\partial\Omega_{2}\subset\partial\mathcal{A}_{3}$. Similar to the proof of lemma 5.1, we can show that $\Omega$ is completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$. ∎
The proof of Theorem 1.3 is similar to the proof of Theorem 1.2 because, after linearizing (1.22) by equation (5.2), (1.22) become
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{2\pi T^{2}(x)}e^{-\frac{v^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (5.23)
Then, equation (5.10) becomes
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (5.24)
$=\frac{1}{\kappa}\int_{\mathbb{R}^{3}}v_{3}hdv+\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(f)dv$
$=\frac{1}{\kappa}\int_{\mathbb{R}^{3}}v_{3}h[f]dv+\frac{1}{\kappa}\int_{\mathbb{R}^{3}}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(h[f])dv+\frac{1}{\kappa}\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(h[f])dv$
$\quad+\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{n+1}\circ\overline{P}(f)dv$
$\coloneq\frac{1}{\kappa}\Big(C_{1}+C_{2}\Big)+D.$
Since the upper bound of $|C_{1}|+|C_{2}|$ can be estimated independently of $n$ and $D$ can be made arbitrary small when $n$ is sufficiently large, the proof is complete.
Chapter 6. Appendix A. The sufficient conditions for unsealed sets
Let $\Omega$ be a subset in $\mathbb{R}^{3}$ and $\partial B$ be a subset of $\partial\Omega$. Determining whether $\Omega$ is unsealed with respect to $\partial B$ is a difficult problem. In this section, we provide some sufficient conditions concerning this problem.
Lemma 6.1.
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{1}$ boundary and $\partial\Omega_{1,a},\partial\Omega_{1,b}$ be subsets of $\partial\Omega$. If $\partial\Omega_{1,a}\subseteq\partial\Omega_{1,b}$ and $\Omega$ is unsealed with respect to $\partial\Omega_{1,a}$, then $\Omega$ is unsealed with respect to $\partial\Omega_{1,b}$.
Proof.
Since $\partial\Omega_{1,a}\subseteq\partial\Omega_{1,b}$, $\big(\partial\Omega\setminus\partial\Omega_{1,b}\big)\subseteq\big(\partial\Omega\setminus\partial\Omega_{1,a}\big)$. Then, by definition 4.1, the proof is complete. ∎
Proposition 6.1.
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with $C^{1}$ boundary ,and let $\partial\Omega_{1}$ be an open subset of $\partial\Omega$. If there exist $\partial C_{1}$, $\partial C_{2}\subset\partial\Omega\setminus\partial\Omega_{1}$ and $0<\zeta<1$ such that
  1. 1.
    $\partial C_{1}\cup\partial C_{2}=\partial\Omega\setminus\partial\Omega_{1}$.
  2. 2.
    $\int_{v\cdot n(x)>0}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(q(x,v))m(v)v\cdot n(x)dv\leq\zeta\quad\forall x\in\partial C_{1}$.
  3. 3.
    $\int_{v\cdot n(x)>0}\textbf{1}_{\partial C_{2}}(q(x,v))m(v)v\cdot n(x)dv=0\quad\forall x\in\partial C_{2}$.
Then, $\Omega$ is unsealed with respect to $\partial\Omega_{1}$.
Proof.
For convenience, we use same symbol in Chapter 4
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{1})d\sigma_{0,1}$
$=\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{1}}(x_{1})d\sigma_{0,1}+\displaystyle\int\textbf{1}_{\partial C_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{2}}(x_{1})d\sigma_{0,1}+\displaystyle\int\textbf{1}_{\partial C_{2}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{2}}(x_{1})d\sigma_{0,1}$
$\leq\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{1}}(x_{1})d\sigma_{0,1}+\displaystyle\int\textbf{1}_{\partial C_{2}}(x_{1})d\sigma_{0,1}.$
If $x\in\partial C_{2}$, then
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{1}}(x_{1})d\sigma_{0,1}\leq\zeta.$
Otherwise,
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{1})d\sigma_{0,1}$ $\leq\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial C_{1}}(x_{1})d\sigma_{0,1}+\displaystyle\int\textbf{1}_{\partial C_{2}}(x_{1})d\sigma_{0,1}$
$\leq\int\displaystyle\textbf{1}_{\partial C_{1}}(x_{1})d\sigma_{0,1}+\displaystyle\int\textbf{1}_{\partial C_{2}}(x_{1})d\sigma_{0,1}$
$=\displaystyle\int_{v\cdot n(x,v)>0}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(q(x,v))m(v)v\cdot n(x)dv$
$\leq\zeta.$
So
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{1})d\sigma_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1}.$ (6.1)
Chapter 7. Appendix B The equivalent definition of solution
Revision In this chapter, we discuss the definition of a solution of (1.5), (1.9) and valid that (5.9) exists when Theorem 1.1 holds. Also, we discuss the relation between the mild solution and $v\cdot\nabla f$.
Definition 7.1.
$F$ is an mild solution of (1.5) if
$F$ $=\mathbf{1}_{\partial\Omega^{-}_{1}}F_{in}(q(x,v),v)+\mathbf{1}_{\partial\Omega^{-}_{2}}m_{T(q(x,v))}\int_{u\cdot n(q(x,v))>0}F(q(x,v),v)u\cdot n(q(x,v))du$
$+\frac{1}{\kappa}\int_{0}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Let
$\nu_{F}(x,v)=\int_{\mathbb{R}^{3}}\int_{S^{2}}\mathbf{B}(|v-u|,\sigma)F(x,v)d\sigma du.$
We have another type of definition to solution of (1.5). That is,
Definition 7.2.
$F$ is an exponential solution of (1.5) if
$F$ $=\mathbf{1}_{\partial\Omega^{-}_{1}}e^{-\int_{0}^{\tau_{-}(x,v)}\frac{1}{\kappa}\nu_{F}(x-sv,v)ds}F_{in}(q(x,v),v)$
$+\mathbf{1}_{\partial\Omega^{-}_{2}}e^{-\int_{0}^{\tau_{-}(x,v)}\frac{1}{\kappa}\nu_{F}(x-sv,v)ds}m_{T(q(x,v))}\int_{u\cdot n(q(x,v))>0}F(q(x,v),v)u\cdot n(q(x,v))du$
$+\frac{1}{\kappa}\int_{0}^{\tau_{-}(x,v)}e^{-\int_{0}^{t}\frac{1}{\kappa}\nu_{F}(x-sv,v)ds}Q_{+}(F,F)(x-tv)dt.$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$. Here,
$Q_{+}(F,G)(x,v)=\int_{\mathbb{R}^{3}}\int_{S^{2}}\left[F(v^{\prime})G(u^{\prime})+F(u^{\prime})G(v^{\prime})\right]B(|v-u|,\omega)d\omega du,$
After linearizing (1.5), we obtain (1.9). Similar (1.5), we can define the mild solution as follow.
Definition 7.3.
$f$ is an mild solution of (1.9) if
$f$ $=\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}(q(x,v),v)$
$+\mathbf{1}_{\partial\Omega^{-}_{2}}\left[m^{-1}m_{T(q(x,v))}P(f)(q(x,v),v)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(q(x,v))}-m)\right]$
$+\frac{1}{\kappa}\int_{0}^{\tau_{-}(x,v)}Q(M+m^{1/2}f,M+m^{1/2}f)(x-tv)dt.$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Also, we can define the exponential solution
Definition 7.4.
$f$ is an exponential solution of (1.9) if
$f=J^{\kappa}\left[f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}\right]+J^{\kappa}\left[m^{-1}m_{T(x)}P(f)+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\textbf{1}_{\partial\Omega^{-}_{2}}\right]+\frac{1}{\kappa}S^{\kappa}_{\Omega}Kf+\frac{1}{\kappa}S^{\kappa}_{\Omega}\Gamma(f,f).$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
It is obvious that the existence of a mild solution to (1.5) is equivalent to the existence of a mild solution to (1.9). Since the proof of the equivalence between Definition 7.1 and Definition 7.2 is similar to that between Definition 7.3 and Definition7.4, we only present the proof for the first case. The statement is as follows.
Theorem 7.1.
Suppose that $\nu_{F}(x-sv,v)$ and $Q_{+}(F,F)(x-sv,v)$ belong to $L^{1}_{loc}(-\tau_{-}(x,-v),\tau_{-}(x,v))$ for almost all $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $F\in L^{1}(\Omega\times\mathbb{R}^{3})$. Then, the definition of 7.1 is equivalent to the definition of 7.2.
Proof.
Let
$g(x,v)=\mathbf{1}_{\partial\Omega^{-}_{1}}F_{in}(q(x,v),v)+\mathbf{1}_{\partial\Omega^{-}_{2}}m_{T(q(x,v))}\int_{u\cdot n(q(x,v))>0}F(q(x,v),v)u\cdot n(q(x,v))du$
and
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt,$
for any $s\in[-\tau_{-}(x,-v),\tau_{-}(x,v)]$. By the definition, $G(0)=F(x,v)$. Moreover, we have
$\int_{\Omega\times\mathbb{R}^{3}}F(x,v)dxdv$ $=\int_{\partial\Omega^{-}}\int_{0}^{\tau_{-}(x,-v)}F(x+sv,v)|v\cdot n(x)|dsdSdv.$
and
$\int_{\Omega\times\mathbb{R}^{3}}G(0)dxdv$ $=\int_{\partial\Omega^{-}}\int_{0}^{\tau_{-}(x,-v)}G(-s)|v\cdot n(x)|dsdSdv.$
Since $G(0)=F(x,v)$, G(s)=F(xsv)G(s)=F(x-sv) for almost all $s\in[-\tau_{-}(x,-v),\tau_{-}(x,v)]$.
$\left\{\begin{aligned} \frac{d}{ds}G(s)&=-\frac{1}{\kappa}Q(F,F)(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)F(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)G(s),&&\text{ in }s\in[-\tau_{-}(x,-v),\tau_{-}(x,v)].\\ G(0)&=F(x,v).\end{aligned}\right$. (7.1)
Note that the solution of (7.1) is unique and $G$ has two representations. One of them is
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt.$
The other is
$G(s)=e^{-\int_{0}^{s}\frac{1}{\kappa}\nu_{F}(x-sv,v)ds}g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}e^{-\int_{0}^{t}\frac{1}{\kappa}\nu_{F}(x-sv,v)ds}Q_{+}(F,F)(x-tv)dt.$
By uniqueness, the proof is complete. ∎
Now, we consider the following boundary value problem.
[…]
We pay attention on the following question
”Does the existence of a mild solution of implies $v\cdot\nabla_{x}F$ exists in distribution sense.”
The mild solution of (7.2) is
$F(x,v)=F_{in}(q(x,v),v)+\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds.$
Let $\varphi$ be a smooth function with compact support. We show the following statement
Theorem 7.2.
If $F$ is a mild solution of (7.2) and $F,G\in L^{p}(\Omega\times\mathbb{R}^{3})$ for some $p\geq 1$, then
$\int_{\Omega\times\mathbb{R}^{3}}F(x,v)v\cdot\nabla_{x}\varphi dxdv=-\int_{\Omega\times\mathbb{R}^{3}}G(x,v)\varphi dxdv$
for any $\varphi\in C^{\infty}_{c}(\Omega\times\mathbb{R}^{3})$.
Proof.
Notice that
$F(x+hv,v)-F(x,v)$ $=\int_{0}^{\tau_{-}(x+hv,v)}G(x+hv-sv,v)ds-\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds$
$=\int_{0}^{\tau_{-}(x,v)+h}G(x-(s-h)v,v)ds-\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds$
$=\int_{-h}^{\tau_{-}(x,v)}G(x-sv,v)ds-\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds$
$=\int_{0}^{h}G(x+sv,v)ds$
for any $h>0$. Let $y=x+hv$. Since $\Omega$ is open and $\varphi$ has a compact support, we yield
$\lim_{h\to 0}\int_{\Omega\times\mathbb{R}^{3}}\frac{1}{h}\int_{0}^{h}G(x+sv,v)ds\varphi dxdv$ $=\lim_{h\to 0}\int_{\Omega\times\mathbb{R}^{3}}\frac{F(x+hv,v)-F(x,v)}{h}\varphi dxdv$
$=\lim_{h\to 0}\int_{\Omega\times\mathbb{R}^{3}}\frac{\varphi(x-hv,v)-\varphi(x,v)}{h}F(x,v)dxdv$
$+\lim_{h\to 0}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\mathbf{1}_{\Omega}(x-hv)-\mathbf{1}_{\Omega}(x)}{h}\varphi(x-hv,v)F(x,v)dxdv$
$=-\int_{\Omega\times\mathbb{R}^{3}}F(x,v)v\cdot\nabla_{x}\varphi dxdv.$
Moreover, we have
$\lim_{h\to 0}\int_{\Omega\times\mathbb{R}^{3}}\frac{1}{h}\int_{0}^{h}G(x+sv,v)ds\varphi dxdv=\int_{\Omega\times\mathbb{R}^{3}}G(x,v)\varphi dxdv.$
Hence the proof is complete. ∎
Chapter 8. The $H^{2}$ estimate of Laplace equation with Robin boundary condition
Revision In this chapter, we prove the inequality (2.12). That is, we discuss about the following elliptic equation
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \hat{\gamma}(x)U+\partial_{\nu}U&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.1)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ but $\hat{\gamma}>0$ on a positive measure subset of $\partial\Omega$. Before we start our argument, we recall several classical trace theorem and elliptic estimate.
Theorem 8.1.
Let $u,v\in W^{1}(\Omega)$ and $\Delta u,\Delta v\in L^{2}(\Omega)$ . Then:
  1. (i)
    $\int_{\Omega}\Delta u\,dx=\int_{\partial\Omega}\partial_{\nu}u\,dS.$
  2. (ii)
    $\int_{\Omega}\nabla v\cdot\nabla u\,dx=-\int_{\Omega}u\Delta v\,dx+\int_{\partial\Omega}u\,\partial_{\nu}v\,dS.$
  3. (iii)
    $\int_{\Omega}\bigl(u\,\Delta v-v\,\Delta u\bigr)\,dx=\int_{\partial\Omega}\left(u\,\partial_{\nu}v-v\,\partial_{\nu}u\right)\,dS.$
Theorem 8.2 ([18] Theorem 1.5.1.3).
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with a Lipschitz boundary $\Gamma$. Then the mapping
$u\mapsto\gamma u$
which is defined for $u\in C^{0,1}(\overline{\Omega})$, has a unique continuous extension as an operator from $W^{1}_{p}(\Omega)$ onto $W^{1-\frac{1}{p}}_{p}(\Gamma)$. This operator has a right continuous inverse independent of $p$.
The following theorem is a general version of the above theorem
Theorem 8.3 ([18] Theorem 1.5.1.2).
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with boundary $\Gamma=\partial\Omega$ of class $C^{k,1}$. Assume that
$s-\frac{1}{p}\notin\mathbb{N},\qquad s\leq k+1,\qquad s-\frac{1}{p}=l+\sigma,$
where $l\in\mathbb{N}_{0}$ and $0<\sigma<1$. Then the mapping
$u\longmapsto\bigl(\gamma u,\,\gamma\partial_{\nu}u,\,\ldots,\,\gamma\partial_{\nu}^{\,l}u\bigr),$
initially defined for $u\in C^{k,1}(\overline{\Omega})$, admits a unique continuous extension as an operator
$W_{p}^{\,s}(\Omega)\;\longrightarrow\;\prod_{j=0}^{l}W_{p}^{\,s-j-\frac{1}{p}}(\Gamma),$
which is onto. Moreover, this operator possesses a continuous right inverse which does not depend on $p$.
Theorem 8.4 ([18] Boundary $H^{2}$-regularity).
Assume
$f\in L^{2}(\Omega).$
Suppose that $u\in H^{1}(U)$ is a weak solution of the elliptic boundary-value problem
$\begin{cases}\Delta u=f&\text{in }\Omega,\\ \gamma\partial_{\nu}u=0&\text{on }\partial\Omega.\end{cases}$
Assume finally
$\partial\Omega\text{ is }C^{2}\text{ and }\int_{\Omega}f=0.$
Then
$u\in H^{2}(\Omega),$
and we have the estimate
$\|u\|_{H^{2}(\Omega)}\leq C\|f\|_{L^{2}(\Omega)},$
the constant $C$ depending only on $\Omega$.
Now, we prove the following estimate
Lemma 8.1.
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Suppose that $U\in H^{1}(\Omega)$ is a weak solution of (8.1). Then
$U\in H^{2}(\Omega),$
and we have the estimate
$\|u\|_{H^{2}(\Omega)}\leq C\left(\|F\|_{L^{2}(\Omega)}+\|u\|_{H^{1}(\Omega)}\right),$
the constant $C$ depending only on $\Omega$.
Proof.
By Green’s identity, we have
$\int_{\Omega}Fdx=\int_{\partial\Omega}\hat{\gamma}(x)UdS.$ (8.2)
We consider the following elliptic problem
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \partial_{n}U&=g\coloneq-\hat{\gamma}(x)U,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.3)
After applying theorem 8.3, there exists $G\in W^{2}(\Omega)$ such that $\partial_{\nu}G=-\hat{\gamma}(x)U$. Then, UGU-G satisfies
$\left\{\begin{aligned} -\Delta(U-G)&=F+\Delta G,&&\text{in }\Omega,\\ \partial_{n}(U-G)&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.4)
By (8.2) and Green’s identity, we have
$\int_{\Omega}F+\Delta Gdx=\int_{\Omega}Fdx+\int_{\Omega}\Delta Gdx=\int_{\Omega}Fdx+\int_{\partial\Omega}\partial_{\nu}GdS=\int_{\Omega}Fdx-\int_{\partial\Omega}\hat{\gamma}(x)UdS=0.$
Applying theorem 8.4, we yield
$\|U-G\|_{H^{2}(\Omega)}\leq C\|F+\Delta G\|_{L^{2}(\Omega)}.$
Using theorem 8.3 and theorem 8.2, we obtain
$\|U\|_{H^{2}(\Omega)}$ $\leq C\left(\|F+\Delta G\|_{L^{2}(\Omega)}+\|G\|_{H^{2}(\Omega)}\right)$
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|G\|_{H^{2}(\Omega)}\right)$
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|\hat{\gamma}(x)U\|_{H^{1/2}(\partial\Omega)}\right)$
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|U\|_{H^{1/2}(\partial\Omega)}\right)$
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|U\|_{H^{1}(\Omega)}\right)$
The proof is complete. ∎
Lemma 8.2.
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Then, (8.1) admits a unique weak solution $U$. Furthermore,
$\|U\|_{H^{1}(\Omega)}\leq C\|F\|_{L^{2}(\Omega)},$ (8.5)
the constant $C$ depending only on $\Omega$.
Proof.
Note that the weak solution of (8.1) means
$\int_{\Omega}\nabla U\cdot\nabla v\,dx+\int_{\partial\Omega}\hat{\gamma}Uv\,dS=\int_{\Omega}Fv\,dx.$ (8.6)
Let
$a(u,v)\coloneq\int_{\Omega}\nabla u\cdot\nabla v\,dx+\int_{\partial\Omega}\hat{\gamma}uv\,dS.$
Since $\hat{\gamma}\in L^{\infty}(\partial\Omega)$ and $u,v\in H^{1}(\Omega)$, we have
$|a(u,v)|\leq C\|u\|_{H^{1}(\Omega)}\|v\|_{H^{1}(\Omega)}.$
Assume that $a$ is not coercive, There exists a sequence in $H^{1}(\Omega)$ such that
$\|u_{n}\|^{2}_{H^{1}(\Omega)}>n\left(\int_{\Omega}|\nabla u_{n}|^{2}\,dx+\int_{\partial\Omega}\hat{\gamma}u_{n}^{2}\,dS\right).$
Without loss of generality, we assume that $\|u_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. We can discover that the weak limit of $\{u_{n}\}$ exists by above inequality. Let $u=w-\lim_{n\to\infty}u$. Observe that
$0=\lim_{n\to\infty}\frac{1}{n}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n\to\infty}\|\nabla u_{n}\|_{L^{2}(\Omega)}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n\to\infty}\int_{\Omega}\nabla u_{n}\cdot\nabla u\,dx=\int_{\Omega}|\nabla u|^{2}\,dx.$
It implies that there exists a subsequence of $u_{n}$ which strongly converge to some $u$ in $H^{1}(\Omega)$. Moreover, we deduce that $u=0$ and obtain a contradiction since $\|u_{n}\|_{L^{2}(\Omega)}=1$. By Lax-Milgram theorem, the proof of the existence is complete. Assume that (8.5) does not hold. There exists a sequence $\{U_{n}\}$ such that
$\|U_{n}\|_{H^{1}(\Omega)}\geq n\|F_{n}\|_{L^{2}(\Omega)}.$
Note that
$\int_{\Omega}|\nabla U_{n}|^{2}\,dx+\int_{\partial\Omega}\hat{\gamma}U_{n}^{2}\,dS=\int_{\Omega}F_{n}U_{n}\,dx\leq\|F_{n}\|_{L^{2}(\Omega)}\|U_{n}\|_{L^{2}(\Omega)}.$
Without loss of generality, we still assume that $\|U_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. After a similar argument, we can immediately conclude that $\{U_{n}\}$ is bounded in $H^{1}(\Omega)$ and there exists a subsequence of $U_{n}$ which strongly converge to 0. It is a contradiction. The uniqueness immediately comes from (8.6) and the similar argument. ∎
After applying the above estimate, we can prove the proposition 2.6 here.
Proof.
The Green’s identity provides the following weak version of (2.8):
$\int_{\partial\Omega\times\mathbb{R}^{3}}\psi f\,dSdv-\iint_{\Omega\times\mathbb{R}^{3}}v\cdot\nabla_{x}\psi\,fdxdv=-\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\frac{1}{\kappa}L(\mathrm{I}-\mathbf{P})fdxdv+\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\phi_{1}dxdv.$ (8.7)
We recall that
$P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}.$
The key of the proof is to choose suitable $H^{1}$ test functions $\psi$ to estimate $\langle f,\Phi_{i}\rangle_{v}$. We choose the test function
$\psi=(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0},$ (8.8)
where $\phi_{\langle f,\Phi_{0}\rangle_{v}}$ is the solution of following elliptic problem
$\left\{\begin{aligned} -\Delta\phi_{\langle f,\Phi_{0}\rangle_{v}}&=\langle f,\Phi_{0}\rangle_{v},&&\text{in }\Omega,\\ \hat{\gamma}(x)\phi_{\langle f,\Phi_{0}\rangle_{v}}+\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.9)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ ( but $\gamma>0$ on a positive measure subset of $\partial\Omega$ ). We bound the right hand side of (2.8) by
$\frac{1}{\kappa}\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\|(\mathrm{I}-\mathbf{P}_{L})f\|_{L^{2}_{0,\gamma/2}}+\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\|\phi_{1}\|_{L^{2}_{0,\gamma/2}}.$
We choose $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ such that
$\int_{\mathbb{R}^{3}}\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)\left(\frac{|v|^{2}}{2}-\frac{3}{2}\right)(v_{i})^{2}\,\mu(v)\,dv=0.$ (8.10)
Moreover, we decompose that
$\begin{aligned} f&=\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2},&&\qquad\text{on }\partial\Omega\times\mathbb{R}^{3},\\ f&=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}+(\mathrm{I}-\mathbf{P}_{L})f,&&\qquad\text{on }\Omega\times\mathbb{R}^{3}.\end{aligned}$ (8.11)
By (8.11), since the $\langle f,\Phi_{4}\rangle_{v}$ contribution vanishes in (2.8) due to our choice of $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ and the $\langle f,\Phi_{1}\rangle_{v}$, $\langle f,\Phi_{2}\rangle_{v}$, and $\langle f,\Phi_{3}\rangle_{v}$ contribution vanish in (2.8) due to the oddness, the left-hand side of (2.8) takes the form of
$\int_{\partial\Omega\times\mathbb{R}^{3}}n\cdot v\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\bigl[\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2}\bigr]$ (8.12)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\,\langle f,\Phi_{0}\rangle_{v}\Phi^{2}_{0}dxdv$ (8.13)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\,(\mathrm{I}-\mathbf{P}_{L})f.$ (8.14)
We make an orthogonal decomposition at the boundary $\partial\Omega_{2}$,
$v=(v\cdot n)n+v-(v\cdot n)n=v_{n}n+v_{\perp}.$
The contribution of $Pf=z(x)m^{\frac{1}{2}}=z(x)\Phi_{0}$ is
$\int_{\partial\Omega_{2}\times\mathbb{R}^{3}}v_{n}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}(x)z(x)\Phi^{2}_{0}dSdv$
$=\int_{\partial\Omega_{2}\times\mathbb{R}^{3}}v_{n}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})\,v_{n}\,z(x)\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}^{2}dSdv+\int_{\partial\Omega_{2}\times\mathbb{R}^{3}}v_{n}(|v|^{2}-\beta_{a})\,v_{\perp}\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\,z(x)\Phi_{0}^{2}dSdv.$
The crucial choice of $\psi$ makes the first term vanish due to the Neumann boundary condition, while the second term also vanishes due to the oddness of $v_{n}v_{\perp}$. Therefore, (8.12) and (8.14) are bounded by
$\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\lVert(I-P_{L})(f)\rVert_{L^{2}_{0,\gamma/2}}$
and
$\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\left(\lVert(I-P)f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}+\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}\right)$
The second term (8.13), for $k\neq i$ vanishes due to the oddness. Hence we only have the $k=i$ contribution:
$\sum_{i=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{a})\,(v_{i})^{2}\,\mu\,\partial_{ii}\phi_{a}\,a.$
Using $\Delta_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}={\langle f,\Phi_{0}\rangle_{v}}$ and $\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(dx)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(dx)}$ we obtain
$\|{\langle f,\Phi_{0}\rangle_{v}}\|_{L^{2}(dx)}\lesssim\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert_{L^{2}_{0,-\gamma/2}}+\lVert(I-P)f\rVert_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{1})}+\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}\end{pmatrix}$.
We therefore deduce the lemma.